### v1

# Ptychoformer: A Physics-Guided Deep Learning Framework for Ptychographic Imaging

Abstract
Ptychographic imaging confront limitations in applying deep learning (DL) for retrieval from diffraction patterns. Conventional neural architectures are optimized for natural images, overlooking the unique physical characteristics of diffraction data, including radial intensity decay and coherent information distributed in concentric rings. In this paper, we present Ptychoformer, a physics-guided DL framework for ptychographic imaging that aligns attention mechanisms and feature extraction with these diffraction physics properties through introducing a dual-branch architecture which accounts for both local and non-local dependencies from the patterns. It consists of a Polar Coordinate Attention (PCA) mechanism that is inspired by the Ewald construction in X-ray crystallography to enhance high-frequency component fidelity. Experimental results demonstrate Ptychoformer’s superior performance across both simulated and real data in preserving fine details and suppressing artifacts. On simulated data, Ptychoformer achieves up to 5.4% higher PSNR and 4.2% higher SSIM for amplitude retrieval compared to existing methods. For real experimental data, it demonstrates up to 12.5% higher PSNR and 31.3% higher SSIM for amplitude retrieval. Notably, Ptychoformer maintains robust performance under limited training data and low overlap ratios, outperforming existing models.
Index Terms:
Ptychography, Physics-Guided Deep Learning, Inverse Imaging Problems, Transformer.
I Introduction
[…]
Figure 1: Diffraction pattern characteristics and diffraction physics. (a): Diffraction patterns exhibit a radial distribution of information, with varying requirements for capturing low and high-frequency information. (b): 2D diffraction patterns can be viewed as projections of the intersection between the Ewald sphere and the crystal onto the detector plane. Sphere’s radius is $1/\lambda$, the wavelength of the incident and diffracted beams. Importantly, the spatial adjacency preference observed in natural images’ feature space also holds in the polar coordinate perspective of diffraction patterns.
However, the widespread application of ptychography faces significant challenges, particularly in computational efficiency. As imaging capabilities advance, data volume grows exponentially, overwhelming conventional algorithms. For instance, processing one second of data from a modern synchrotron source (10-megapixel detector, 32-bit depth, 2 kHz, 640 Gb/s) can take up to an hour [7]. Additionally, achieving high-quality retrievals requires 60-70% probe position overlap [8], further increasing computational complexity. These factors severely constrain ptychography’s viability in real-time and high-throughput applications.
To address these limitations, researchers have redefined ptychography as a data-driven supervised learning task, leveraging deep learning (DL) techniques. Notable examples including PtychoNN [9], Deep-phase-imaging (DPI) [10], and PtyNet [11] have demonstrated improved retrieval efficiency. Unlike conventional approaches that require repeated iterations across the spatial and frequency domains, these methods utilize experimentally acquired diffraction patterns and complex amplitude images reconstructed by traditional algorithms as training data. Through convolutional neural network (CNN) based architectures, they establish a direct mapping from diffraction patterns to complex amplitude retrieval, significantly enhancing computational efficiency. Despite their success, CNN-based methods face challenges in processing multi-frequency information, capturing long-range correlations, and distinguishing high-frequency signals from noise in diffraction patterns, as illustrated in Fig. 1(a). These limitations are due to the sliding window operation of convolution kernels in spatial coordinates and the information flattening in multi-layer convolution during transmission [12, 13, 14]. Transformer models, while powerful in various computer vision tasks, have limited application in ptychography due to the mismatch between their design principles and the radial nature of diffraction patterns. For instance, Vision Transformers (ViT), are primarily optimized for natural images in real space (e.g., patch extraction operations and local feature preference priors of the Softmax function). However, these priors do not directly map to the radial nature of diffraction data in reciprocal space shown in Fig. 1(a).
Addressing the inherent limitations of existing methods requires a fundamental reconsideration of deep learning architectures for ptychographic retrieval, with a focus on optimizing neural networks for representing diffraction patterns in reciprocal space. This insight guided the development of Ptychoformer, a novel framework that bridges frequency-domain diffraction patterns and real-space images using physics-informed deep learning techniques. Drawing inspiration from the Ewald construction in X-ray crystallography, as depicted in Fig. 1(b), treating 2D diffraction patterns as projections of the intersections between the Ewald sphere and the reciprocal lattice onto the detector plane, as shown in the left of Fig. 1. This insight revealed that diffraction patterns are inherently defined in spherical rather than planar geometry. Consequently, we developed Polar Coordinate Attention (PCA), an attention mechanism leveraging polar coordinates that naturally aligns with diffraction physics. This polar coordinate-based attention mechanism reframes real-space structural priors in the diffraction context, mapping scattering vector magnitude to $r$ and angular information to $\theta$, thus capturing both radial intensity attenuation and angular coherence. Recognizing the multi-scale and multi-frequency nature of diffraction patterns, as shown in the right side of Fig. 1, Ptychoformer employs a dual-branch architecture combining a Local Dependencies Branch constructed from standard ViT blocks with a NonLocal Coherence block containing our designed PCA. This design balances the capture of long-range and local dependencies in diffraction patterns. Ptychoformer’s architecture intrinsically aligns with ptychography physics, creating a more natural correspondence between model operations and underlying physical processes.
This structural redesign based on physical insights enables our model to more effectively extract and utilize global coherent information and long-range dependencies in diffraction data, thus offering a physics-informed approach to ptychographic retrieval. Our comprehensive experiments on both simulated and real datasets demonstrate Ptychoformer’s significant advantages over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency. Under challenging conditions including low overlap ratios, and limited training data, our model maintains superior retrieval quality, achieving up to 36.1% reduction in Mean Squared Error (MSE) and 12.03% increase in Structural Similarity Index Measure (SSIM) compared to the best-performing existing end-to-end methods. These improvements could significantly enhance the applicability of ptychography in time-sensitive or radiation-sensitive imaging scenarios across various scientific disciplines.
The primary contributions of this work are summarized as follows:
  • We present Ptychoformer, an end-to-end deep learning framework addressing limitations of existing DL-based ptychography methods. Unlike conventional neural architectures optimized for spatial-domain natural images, our approach incorporates designs aligned with diffraction physics.
  • We propose the PCA mechanism, a physics-inspired approach that decomposes attention into radial and angular components, enabling efficient extraction of coherent information across frequency bands in diffraction patterns.
  • Extensive experiments on both simulated and real datasets show that the proposed method achieves high-fidelity and superior artifact reduction, while maintaining efficiency under challenging conditions including noisy environments and limited data acquisition scenarios.
II Related Works
II-A Deep Learning-Based Ptychographic Imaging
Deep learning in ptychographic imaging has recently enhanced computational efficiency and retrieval quality, categorized into three strategies:
II-A1 Pre-processing
Integrating deep learning with iterative algorithms improves initial estimates. The physics-informed automatic differentiation ptychography (ADP) framework uses pre-trained autoencoders to map high-dimensional image data to a low-dimensional latent space [15], while the double deep image prior (DDIP) method reduces the optimization parameter space [16], both enhancing convergence rates and noise robustness.
II-A2 Post-processing
Neural networks refine reconstructions from traditional algorithms, including enhancing a single iteration of the Gerchberg-Saxton-Fienup (GSF) algorithm to improve spatial resolution and reduce artifacts [5], showing promise in live cell imaging.
II-A3 End-to-End
End-to-end methods directly map diffraction patterns to complex object functions, bypassing iterative processes. Examples include PtychoNN with a modified U-Net and two-branch decoder for amplitude and phase [9], PtyNet with group convolution and Leaky ReLU for efficiency [11], and DPI using a traditional U-Net with skip connections [10].
Despite improved retrieval speed, these methods face challenges to preserve high-frequency information and the global coherence of diffraction patterns, as CNNs and standard Vision Transformers (ViTs) are tailored for real-space images with local correlations. Recent studies using ViTs for ptychographic retrieval leverage self-attention for long-range dependencies [17], but applying ViT architectures to frequency-domain data is limited by radial intensity decay and concentric information distribution. Addressing these frequency-domain challenges, including processing global coherence and preserving high frequencies, remains crucial for advancing retrieval quality and efficiency, especially with limited data or low overlap ratios.
II-B ViT-based Methods
Figure 2: The proposed Ptychoformer architecture for ptychographic imaging. It features dual branches: a Local Dependencies Branch with standard ViT blocks, and a NonLocal Coherence Branch with PolarCoordinate Attention mechanism. The model processes logarithmically mapped diffraction patterns, combining features from both branches before decoding into amplitude and phase images with physical constraints applied.
Diffraction patterns, governed by wave optics, approximate the Fourier transform of an object’s transmission function in far-field conditions [18], resulting in reciprocal space features like sparse representations and concentric rings [19]. Accurate image retrieval and phase retrieval require effectively capturing both sparsity and long-range dependencies within these patterns. Traditional CNNs, though effective for natural images [20, 21], are limited with the global nature of diffraction phenomena due to their local receptive fields. Vision Transformers (ViTs), introduced by Dosovitskiy et al. [22], offer a promising alternative with self-attention mechanisms that capture global features and long-range dependencies [23], outperforming CNNs in various computer vision tasks [24, 25] and extending to fields like medical image segmentation [24, 26], climate modeling [27], and molecular structure prediction [28]. However, applying ViTs to phase retrieval in CDI is challenging due to the lack of CNN-like inductive biases such as translational invariance and local connectivity, coupled with often insufficient data in practical phase imaging. For example, PtychoDV [17] uses ViTs for feature extraction from diffraction patterns but shows inferior retrieval performance compared to CNN-based U-Net architectures without iterative post-processing. This gap highlights the potential for developing hybrid methods that leverage transformers’ global feature extraction while addressing data limitations in ptychography.
III Methodology
III-A Problem Formulation
In ptychography, we aim to reconstruct a complex-valued object function $O(\mathbf{r})$, where $\mathbf{r}$ is the position vector in real space. This retrieval is based on a set of diffraction patterns measured at spatial frequencies $\mathbf{q}$ in reciprocal space. The process involves a probe function $P(\mathbf{r})$, which interacts with the object at various scanning positions $\{\mathbf{r}_{j}\}_{j=1}^{J}$. The ptychography problem can be formulated as:
$\mathcal{F}\{\psi_{j}(\mathbf{r})\}=\sqrt{I_{j}(\mathbf{q})}\cdot\exp(i\phi_{j}(\mathbf{q}))$ (1)
where $\psi_{j}(\mathbf{r})=P(\mathbf{r}-\mathbf{r}_{j})\cdot O(\mathbf{r})$ is the exit wave function, $\mathcal{F}$ denotes the Fourier transform, $I_{j}(\mathbf{q})$ is the measured diffraction intensity, and $\phi_{j}(\mathbf{q})$ is the phase of the diffraction pattern.
To account for real-world factors, we model the diffraction intensity as:
$\begin{split}I_{j}(\mathbf{q})=&\mathcal{P}_{\eta}\left(\eta\left|\mathcal{F}\left\{\mu P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})O(\mathbf{r})\right.\right.\right.\\ &\left.\left.\left.+(1-\mu)P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})\mathbb{E}_{\mathbf{r}}[O(\mathbf{r})]\right\}\right|^{2}\right)\\ &+\mathcal{N}(0,\sigma^{2})\end{split}$ (2)
where $\mu$ is the coherence parameter, $\eta$ is the detector efficiency, $\mathcal{P}_{\eta}$ models Poisson shot noise, $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise, and $\delta\mathbf{r}_{j}$ accounts for positional jitter. The objective function for ptychographic retrieval is defined as:
$\begin{split}\mathcal{J}(O,P)=&\frac{1}{J}\sum_{j=1}^{J}\left\|I_{j}-|\mathcal{F}\{P(\mathbf{r}-\mathbf{r}_{j})\cdot O(\mathbf{r})\}|^{2}\right\|_{2}^{2}\\ &+\lambda_{1}\Omega_{1}(O)+\lambda_{2}\Omega_{2}(P)\end{split}$ (3)
where $\left\|\mathbf{x}\right\|_{2}^{2}=\sum_{i}x_{i}^{2}$ is the squared Euclidean norm, $\Omega_{1}(O)=\left\|\nabla O\right\|_{1}$ enforces sparsity in the object gradient, $\Omega_{2}(P)=\left\|P-P_{0}\right\|_{2}^{2}$ constrains the probe function to maintain proximity to to an initial estimate $P_{0}$, and $\lambda_{1}$ and $\lambda_{2}$ are regularization weights.
III-B Ptychoformer Architecture
Figure 3: Architectural overview of the PCA mechanism. The diagram illustrates the flow of input features through polar coordinate transformation, query-key-value attention computations, and subsequent operations including polar coordinate weighting and Hadamard product. The process culminates in multi-head attention, followed by concatenation and linear projection to produce the final output.
The Ptychoformer architecture uses a bifurcated structure to map diffraction patterns to a reconstructed complex-valued object function. The input undergoes a logarithmic transformation $I_{\text{log}}(\mathbf{q},j)=\log(1+I_{j}(\mathbf{q}))$ as a preprocessing step to address the large disparity between high and low frequency values in the diffraction patterns.
III-B1 Local Dependencies Branch
The Local Dependencies Branch employs a standard ViT to analyze diffraction patterns, aiming to capture local features and relationships within the pattern. The branch’s structure consists of several key components. Initially, a patch extraction mechanism segments the diffraction pattern into smaller sections. These segments then undergo reshaping to facilitate subsequent processing. Following this, an encoding step is applied to each patch, designed to emphasize relevant features. The encoded patches are then processed through a series of ViT blocks. Each ViT block $\mathcal{B}_{l}$ in $\mathcal{B}_{L}$ is defined as:
$\begin{split}\mathcal{B}_{l}(\mathbf{X})&=\mathbf{X}+\mathcal{FFN}(\mathcal{LN}(\mathbf{X}+\mathcal{MSA}(\mathcal{LN}(\mathbf{X}))))\\ \mathcal{MSA}(\mathbf{X})&=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})\mathcal{W}^{O}\\ \text{head}_{i}&=\text{Attention}(\mathbf{X}\mathcal{W}^{Q}_{i},\mathbf{X}\mathcal{W}^{K}_{i},\mathbf{X}\mathcal{W}^{V}_{i})\end{split}$ (4)
where $\mathcal{LN}$ denotes layer normalization, $\mathcal{MSA}$ is multi-head self-attention with $h$ heads, $\mathcal{FFN}$ is a position-wise feed-forward network, and $\mathcal{W}^{Q}_{i},\mathcal{W}^{K}_{i},\mathcal{W}^{V}_{i},\mathcal{W}^{O}$ are learnable weight matrices. Despite apparent disruption, patch extraction facilitates low-frequency information capture, especially from central patches containing crucial low-frequency components.
III-B2 NonLocal Coherence Branch
The NonLocal Coherence Branch operates directly on the full-resolution input, intentionally omitting patch extraction to preserve the global structure of the diffraction pattern. It replaces the multi-head self-attention in standard ViT blocks with PCA, and the radial information introduced by PCA eliminates the need for position embeddings.
While traditional self-attention mechanisms are effective for capturing long-range dependencies, they are not optimized for ptychographic data. Designed for natural images, they incorporate spatial positional encodings $\text{PE}(p)$ that can bias attention towards local regions: $A_{pq}=\text{softmax}(\frac{(x_{p}+\text{PE}(p))^{T}(x_{q}+\text{PE}(q))}{\sqrt{d_{k}}})$. This provides structural priors suited for conventional visual tasks but may not be optimal for frequency-domain diffraction patterns. To address this limitation, we introduce the PCA mechanism. PCA transforms spatial coordinates into polar coordinates, shifting the structural priors from Cartesian spatial locality to the concentric information distribution inherent in frequency-domain diffraction patterns. This transformation enables PCA to better capture the radial and angular dependencies characteristic of ptychographic data. The PCA layer $\mathcal{PCA}:\mathbb{R}^{B\times H\times W\times C}\to\mathbb{R}^{B\times H\times W\times C^{\prime}}$ is defined as:
$\mathcal{PCA}(\mathbf{X})=\text{Reshape}(\mathcal{F}(\text{Flatten}(\mathbf{X})))$ (5)
$\mathcal{F}(\mathbf{X})=\text{Concat}(\text{Head}_{1},\text{Head}_{2},\dots,\text{Head}_{h})\mathcal{W}_{O}$ (6)
$\text{Head}_{i}=\text{softmax}(\mathcal{A}r^{i}(\mathbf{X})+\mathcal{A}\theta^{i}(\mathbf{X}))\mathcal{V}^{i}(\mathbf{X})$ (7)
$\mathcal{A}r^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{R}$ (8)
$\mathcal{A}\theta^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{\Theta}$ (9)
where $\mathcal{Q}^{i},\mathcal{K}^{i}\in\mathbb{R}^{N\times d_{k}}$, $\mathcal{V}^{i}\in\mathbb{R}^{N\times d_{v}}$ are the linear projections for queries, keys, and values for the $i$-th head, $d_{k}$ is the dimension of the keys, $\mathbf{R},\mathbf{\Theta}\in\mathbb{R}^{N\times N}$ are the normalized radial and angular coordinate matrices, respectively, $\odot$ denotes the Hadamard product (element-wise multiplication), and $\mathcal{W}_{O}\in\mathbb{R}^{hd_{v}\times C^{\prime}}$ is the output projection matrix. Here, $N=H\times W$ is the number of spatial positions. The coordinate matrices are computed as:
$R_{ij}=\frac{\sqrt{(i-W/2)^{2}+(j-H/2)^{2}}}{\sqrt{(W/2)^{2}+(H/2)^{2}}}$ (10)
$\Theta_{ij}=\frac{\arctan 2(j-H/2,i-W/2)}{\pi}$ (11)
where $i\in[1,W]$ and $j\in[1,H]$ are the horizontal and vertical indices of the spatial positions, respectively. This mechanism decomposes attention into radial and angular components, mirroring the symmetry of diffraction patterns. Operating without patch extraction, this branch preserves global structure. Its polar coordinate attention is designed to extract global coherence information, particularly high-frequency components requiring long-range dependency modeling.
III-B3 Feature Fusion and Decoding
Let $\mathcal{F}_{\text{Local}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{NonLocal}}$ denote the feature extraction operators from the Local Dependencies Branch and the NonLocal Coherence Branch, respectively. The decoder is defined as a composition of three upsampling operators $\mathcal{D}=\mathcal{D}_{3}\circ\mathcal{D}_{2}\circ\mathcal{D}_{1}$, each progressively increasing the spatial resolution using transposed convolutions. Here, $\circ$ denotes function composition, such that $(f\circ g)(x)=f(g(x))$ for any compatible functions $f$ and $g$. The feature fusion process, combining features from both branches, is formulated as:
$\mathcal{F}_{\text{fused}}=\mathcal{C}(\mathcal{F}_{\text{Local}},\mathcal{F}_{\text{NonLocal}})$ (12)
where $\mathcal{C}$ concatenates $\mathcal{F}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}$, followed by a convolution to integrate the features. The fused features $\mathcal{F}_{\text{fused}}$ are then processed through the three-layer decoder $\mathcal{D}$, implemented using transposed convolutions with batch normalization and ReLU activations, split into parallel branches $\mathcal{D}_{\text{amplitude}}$ and $\mathcal{D}_{\text{phase}}$, to reconstruct the preliminary amplitude $\mathbf{A}(\mathbf{r})$ and phase $\boldsymbol{\phi}(\mathbf{r})$ images.
III-B4 Output Constraints
To ensure the physical validity of the reconstructed complex-valued object function, we apply constraints to the final outputs. The amplitude $\mathbf{A}_{\text{final}}:\mathbb{R}^{2}\to\mathbb{R}^{+}$ and phase $\boldsymbol{\phi}_{\text{final}}:\mathbb{R}^{2}\to[-\pi,\pi]$ are constrained as:
$\mathbf{A}_{\text{final}}(\mathbf{r})=A_{\text{max}}\cdot\sigma\left(\frac{\mathbf{A}(\mathbf{r})}{A_{\text{max}}}\right),\quad\boldsymbol{\phi}_{\text{final}}(\mathbf{r})=\pi\cdot\tanh\left(\boldsymbol{\phi}(\mathbf{r})\right)$ (13)
where $A_{\text{max}}$ is the maximum allowable amplitude, $\sigma$ is the sigmoid function, and $\tanh$ is the hyperbolic tangent function. The final complex-valued object function is reconstructed as
$O_{\text{reconstructed}}(\mathbf{r})=\mathbf{A}_{\text{final}}(\mathbf{r})\cdot\exp\left(i\boldsymbol{\phi}_{\text{final}}(\mathbf{r})\right).$
III-B5 Training Objective
The loss function $\mathcal{L}:\Theta\to\mathbb{R}^{+}$ is defined as:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{n=1}^{N}\left[\left\|A_{\text{true}}^{(n)}-A_{\text{pred}}^{(n)}\right\|_{2}^{2}+\left\|\phi_{\text{true}}^{(n)}-\phi_{\text{pred}}^{(n)}\right\|_{2}^{2}\right]$ (14)
where $N$ is the number of training samples, $A_{\text{true}}^{(n)}$ and $\phi_{\text{true}}^{(n)}$ are the true amplitude and phase for the $n$-th sample, respectively, and $A_{\text{pred}}^{(n)}=\mathbf{A}_{\text{final}}^{(n)}$ and $\phi_{\text{pred}}^{(n)}=\boldsymbol{\phi}_{\text{final}}^{(n)}$ are the corresponding model predictions after applying the output constraints. The optimization problem is formulated as:
$\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)$ (15)
where $\theta^{*}$ represents the optimal model parameters that minimize the loss function $\mathcal{L}(\theta)$ over the training dataset.
IV Experiments and Results
IV-A Experiment Setup
IV-A1 Implementation Details
The model was trained using a batch size of 32 for 25 epochs. Adam optimizer with ReduceLROnPlateau callback (factor 0.5, patience 2, $\min_{\operatorname{lr}}$ 0.0001) adjusts learning rate based on validation loss. Training on Tesla T4 GPU. Custom callback monitors memory, time, and convergence. Model compiled with MSE loss. ReduceLROnPlateau strategy implemented. Dataset shuffled with 5% validation split.
IV-A2 Datasets
We utilize two types of datasets in our experiments: Simulated Dataset: Dataset based on magnetic nanostructure simulation [29] using Ni80Fe20 parameters. 512×512 LTEM images represent 5×5 nm areas. The dataset contains 2982 images with 75% overlap and 3-pixel jitter. Gaussian noise simulates realism, with added Poisson and Gaussian noise for readout and photon effects. Amplitude and phase undergo minmax normalization. Real Experimental Data: Experimental data from Argonne National Laboratory [9] comprises 16,100 triplets of diffraction data, amplitude, and phase images. 161 × 161 point scan at 30 nm increments from X-ray nanoprobe beamline 26-ID. PIE algorithm generates ground truth. For both datasets, we split the data into 80% for training and 20% for testing. The validation set is created using 5% of the training data.
IV-A3 Evaluation Metrics
The performance of the models is evaluated using MSE, Peak Signal-to-Noise Ratio (PSNR), and SSIM, with the ground truth values of the materials basis images as references.
Figure 4: Performance comparison of single-shot experiment results and full-stitched scene retrieval. (a) and (b) show two representative scan points. (c) represents the full-stitched scene retrieval across models.
TABLE I: Quantitative comparison of full-stitched scenes under noise-free and noisy conditions. Results are presented as mean $\pm$ standard deviation from ten independent experiments.
Condition Method Amplitude Phase
MSE (×$\times 10^{-2}$) $\downarrow$ PSNR (dB) $\uparrow$ SSIM (%) $\uparrow$ MSE (×$\times 10^{-2}$) $\downarrow$ PSNR (dB) $\uparrow$ SSIM (%) $\uparrow$
Noise-free PtychoNN 4.32 $\pm$ 0.02 13.52 $\pm$ 0.17 83.50 $\pm$ 0.80 1.22 $\pm$ 0.01 19.09 $\pm$ 0.19 67.10 $\pm$ 0.70
PtyNet 4.46 $\pm$ 0.03 13.56 $\pm$ 0.15 83.10 $\pm$ 0.60 1.28 $\pm$ 0.01 19.10 $\pm$ 0.26 67.80 $\pm$ 1.30
DPI 4.29 $\pm$ 0.03 13.68 $\pm$ 0.12 83.40 $\pm$ 0.50 1.26 $\pm$ 0.01 18.91 $\pm$ 0.24 67.70 $\pm$ 1.10
Ptychoformer 3.63 $\pm$ 0.04 14.42 $\pm$ 0.08 87.00 $\pm$ 0.30 1.14 $\pm$ 0.01 19.41 $\pm$ 0.05 70.60 $\pm$ 0.70
Noisy PtychoNN 6.66 $\pm$ 0.03 11.76 $\pm$ 0.12 74.00 $\pm$ 0.89 1.47 $\pm$ 0.04 18.33 $\pm$ 0.17 58.90 $\pm$ 1.00
PtyNet 6.50 $\pm$ 0.02 11.91 $\pm$ 0.13 74.80 $\pm$ 0.66 1.43 $\pm$ 0.07 18.46 $\pm$ 0.18 62.10 $\pm$ 1.50
DPI 6.46 $\pm$ 0.02 11.76 $\pm$ 0.10 74.20 $\pm$ 0.65 1.48 $\pm$ 0.06 18.31 $\pm$ 0.20 62.00 $\pm$ 1.20
Ptychoformer 5.91 $\pm$ 0.03 12.34 $\pm$ 0.11 78.80 $\pm$ 0.59 1.31 $\pm$ 0.06 18.76 $\pm$ 0.16 68.30 $\pm$ 1.10
Note: For noisy conditions, noise is simulated using Gaussian and Poisson distributions to model readout noise and photon noise, respectively. Bold values indicate the best performance for each metric. PSNR in dB, and SSIM in percentage. $\uparrow$: higher is better, $\downarrow$: lower is better.
IV-B Performance Comparison and Analysis
We evaluated Ptychoformer against three mainstreamed CNN-based methods (PtychoNN, PtyNet, and DPI). Our analysis covered single-shot retrieval and full-stitched field retrieval under both ideal and noisy conditions.
IV-B1 Single Scan Point Retrieval Analysis
Ptychoformer produced retrievals with improved edge definition and clarity compared to those from the other methods, particularly evident in the vertical edge on the left side of Fig. 4(a). The reconstructed images also showed contrast levels closer to the ground truth, with more natural transitions between dark and bright areas. By contrast, the CNN-based models tended to smooth these sharp features and introducing artifacts and distortions.
IV-B2 Full-stitched Field Retrieval Analysis
Table I presents full-field stitching results under ideal (noise-free) and realistic (noisy) conditions, demonstrating Ptychoformer’s superior performance across all metrics. In ideal conditions, Ptychoformer achieved 16.0% lower MSEamp and 6.7% higher PSNRamp compared to PtychoNN, with similar improvements under noisy conditions. Fig. 4(c) visually confirms Ptychoformer’s superior global consistency and minimal boundary artifacts, particularly in reconstructing vertical column ends and background textures. The dual-branch structure enables our model to effectively capture both low-frequency and mid-to-high frequency information of diffraction patterns. This capability stems from the model’s ability to capture the intrinsic geometric structure of the data, analogous to continuous mapping on high-dimensional manifolds.
Repeated measures ANOVA and pairwise $t$-tests (Bonferroni correction, $\alpha=0.0083$) showed Ptychoformer significantly outperformed the other methods in all metrics in both conditions ($p<0.0083$), especially in MSEamp and SSIMamp ($p<0.0001$). The three CNN-based methods showed no significant differences, performing similarly across most metrics.
IV-C Validation on Experimental Synchrotron Data
Figure 5: Performance comparison on real experimental samples. (a,b) Visual comparison of retrieved amplitude and phase. For a specific hook-shaped detail, only our model effectively restores it, significantly outperforming CNN-based methods in fine structure retrieval. (c-e) Quantitative comparison of PSNR, SSIM, and MSE.
Fig. 5 demonstrates Ptychoformer’s superior performance on real experimental data, consistent with our findings from simulated datasets. Across all metrics (PSNR, SSIM, and MSE) for both amplitude and phase retrieval, Ptychoformer consistently outperforms other models. These improvements are visible in fine structural details (Fig. 5a,b), indicating Ptychoformer’s enhanced resolution and fidelity in practical imaging applications.
IV-D Generalization Capabilities Comparison
We evaluated Ptychoformer’s generalization by trained on simulated tungsten patterns with straight-line features and evaluated on complex, curved patterns without fine-tuning, assessing the model’s ability to extrapolate learned features to significantly different sample geometries. Fig. 6 shows Ptychoformer’s superior retrieval of curved patterns. Quantitative results in Table II support these visual observations. For amplitude retrieval, Ptychoformer achieves a PSNRamp of 8.92 dB and SSIMamp of 59.06%, outperforming other models by 0.83-0.95 dB and 8.51-11.28%, respectively. For phase retrieval, Ptychoformer attains a PSNRphase of 8.68 dB and SSIMphase of 49.44%, surpassing other models by 0.91-1.04 dB and 32.16-35.71%, respectively. These results demonstrate Ptychoformer’s ability to capture fundamental relationships between diffraction patterns and sample structures, aligning with the geometry-agnostic nature of ptychographic phase retrieval.
Figure 6: Comparison of generalization capabilities across different models: Direct test on a new unseen sample with significantly different distribution from the training set. (a) Amplitude and (b) phase retrieval of a single diffraction pattern. (c) Amplitude and (d) phase retrieval of a full scene (Fudan University logo). Models compared: Ground truth, PtychoNN, PtyNet, DPI, and Ptychoformer.
TABLE II: Performance Comparison for Out-of-Distribution Generalization.
Model Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
PtychoNN 8.03 50.55 7.77 37.34
PtyNet 7.97 47.78 7.64 36.43
DPI 8.09 50.14 7.65 37.41
Ptychoformer 8.92 59.06 8.68 49.44
IV-E Performance Comparison Under Limited Data
Despite significant advancements in ptychography, a major challenge remains: extended data acquisition periods. While higher overlap between scanning points typically yields better retrieval results [14], it also increases experiment duration and radiation exposure. This is particularly problematic for radiation-sensitive materials [30] and in situ dynamic studies [31]. Our study focuses on this aspect, evaluating model performance under two kinds of limited data conditions.
IV-E1 Impact of Training Set Size
Figure 7: Training and testing set divisions of different sizes. Black boxes represent training set sizes, while white boxes represent testing sets, shown as squares.
Fig. 8 (a-c) illustrates Ptychoformer’s performance under varying training set sizes. As the training size decreases from 2928 to 50, Ptychoformer consistently outperforms CNN-based models, especially at smaller training set sizes. Comparing with DPI, the best-performing CNN method, we showcase results at representative training sizes of 1000 and 50. Even with 1000 samples, DPI exhibits limitations with complex structures like the top-right crosshatch pattern, producing block-shaped high-frequency artifacts. Ptychoformer, in contrast, maintains superior resolution. At the minimal training size of 50 samples, DPI’s retrieval quality significantly deteriorates, while Ptychoformer still distinguishes basic structural features like crosshatches and L-shaped patterns. This robustness is attributed to Ptychoformer’s PCA mechanism and dual-branch architecture, which align with diffraction physics and capture multi-scale features effectively, demonstrating superior ability to learn from limited data.
IV-E2 Performance Across Overlap Ratios
Fig. 8 (d) demonstrates Ptychoformer’s effectiveness at varying overlap ratios. At 0% overlap, Ptychoformer achieves higher PSNRamp, surpassing CNNs. From 75% to 0% overlap, Ptychoformer’s PSNRamp drops by a smaller amount compared to CNNs. This resilience, analyzable through compressed sensing theory. [32], suggests Ptychoformer learns an optimal sparse prior. Lower overlap requirements could enable faster experiments with reduced radiation damage. Ptychoformer’s efficiency demonstrates the synergy between deep learning and domain-specific knowledge in ptychography.
Figure 8: Performance comparison across training set sizes and overlap ratios. (a,b) Retrievals with 1000 and 50 training samples. (c,d) PSNR and SSIM metrics for varying training sizes and overlap ratios. Higher values for both PSNR and SSIM indicate better performance.
TABLE III: Ablation Study Results
Model Variant Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
Ours w/o NLB 14.10 86.60 18.39 65.40
Ours w/ MinN 13.82 83.90 18.64 64.10
Ours w/o D 11.02 60.50 17.32 54.50
Ours w/o OC 14.21 86.10 18.39 69.40
Ours 14.42 87.00 19.41 70.60
IV-F Ablation Study
To justify each component of the proposed Ptychoformer, we conducted comprehensive ablation studies. We compared several variants of our model: 1) Ours without NonLocal Coherence Branch (Ours w/o NLB), using only the Local Dependencies Branch; 2) Ours with Multi-Head Self-Attention (MHSA) in NLB (Ours w/ MinN), replacing PCA with standard MHSA in the NonLocal Branch; 3) Ours without Decoder (Ours w/o D), using fully connected layers instead of CNN for upsampling in the decoder; 4) Ours without Output Constraints (Ours w/o OC), removing the amplitude and phase constraints in the output layer; and 5) Ours (Full Model), the complete proposed Ptychoformer model. Table III presents the quantitative results of these variants, with the best performance in bold. The NonLocal Coherence Branch, reflecting diffraction patterns’ global nature, showed a modest impact, primarily affecting structural integrity. PCA demonstrated clear advantages over standard MHSA, with its replacement by MHSA causing a 4.16% drop in PSNRamp and 3.56% in SSIMamp.
The CNN decoder proved crucial, its absence leading to a substantial 23.58% decrease in PSNRamp and 30.46% in SSIMamp. This significant impact is attributed not only to its role in maintaining spatial coherence during upsampling but also to its inductive bias, which enables effective learning from limited data. Output constraints, while showing a modest quantitative impact, play a vital role in enforcing physical consistency. These results validate the effectiveness of each proposed component in our architecture.
V Discussion
V-A Frequency Analysis
Further analysis of the noise-free full-stitched field using 1D diagonal cross-sections of the average 2D power spectral density (PSD) and frequency distribution statistics (Fig. 9). In Fig. 9(a), Ptychoformer’s PSD curve closely follows the ground truth, while the DPI model exhibits significant mid and high-frequency fluctuations, indicating potential artifacts. Quantitatively, Ptychoformer achieves a frequency distribution (99.75%, 0.22%, 0.03% for low, mid, and high frequencies, respectively), closely approximating the ground truth (99.58%, 0.38%, 0.04%). This performance exceeds that of CNN-based models, which show mid and high frequency degradation (e.g., PtychoNN: 99.93%, 0.06%, 0.01%). These results highlight Ptychoformer’s efficacy in extracting mid-to-high frequency information and retrieving corresponding detailed features, leading to superior artifact suppression.
Figure 9: Frequency analysis comparison of full-scene ptychographic retrievals. (a) 1D diagonal cross-sections of average 2D PSD from stitched simulations. Red curve: ground truth; Blue curve: models. The red box represents obvious abnormal model retrieval. (b) Quantitative breakdown of energy distribution across low, mid, and high frequency bands. Frequency ranges are defined based on the radial distance from the PSD center, with boundaries at 1/3 and 2/3 of the maximum frequency.
Figure 10: Performance analysis of weighted loss function $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ in ptychographic retrieval. PSNR, SSIM are plotted against $\alpha$ for amplitude and phase retrievals.
V-B Loss Function Analysis
To optimize retrieval quality in ptychographic phase retrieval, we evaluated various loss functions (as shown in Table IV. Standard metrics like MSE and Mean Absolute Error (MAE) showed similar performance, while the Huber loss [33], defined as: $L_{\text{Huber}}(x)=\begin{cases}\frac{1}{2}x^{2}&\text{for }|x|\leq\delta\\ \delta(|x|-\frac{1}{2}\delta)&\text{otherwise}\end{cases}$ offered robustness to outliers. The Negative Pearson Correlation Coefficient (NPCC) [34] loss, despite its success in single-frame phase retrieval [12], showed poor convergence stability in our experiments. This is likely due to its sensitivity to local statistics in small ptychographic patches, contrasting with its effectiveness on larger, more statistically stable single-frame images.
We introduced a novel combined loss function inspired by the work on image restoration with neural networks [35]: $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ where $L_{\text{NSSIM}}=-\text{SSIM}(I_{q},|\mathcal{F}{P(r)\psi(r)}|^{2})$, where $\psi(r)$ defined as $\psi(r)=P(r-r_{j})O(r)$. This approach balances global consistency (MSE) with local structural preservation ( Negative Structural Similarity Index Measure (NSSIM) [36]). Experiments revealed optimal performance at $\alpha=0.9$, with consistent improvement as $\alpha$ increased from 0.2 to 0.9, as shown in Fig. 10. This robustness to $\alpha$ values reduces the need for precise tuning in practical applications. The combined loss function excels in addressing affine ambiguity, as $\nabla L_{\text{MSE}}$ is sensitive to global affine transformations $T$ ($\frac{\partial L_{\text{MSE}}}{\partial T}\neq 0$), while $\nabla L_{\text{NSSIM}}$ maintains local structure invariance ($\frac{\partial L_{\text{NSSIM}}}{\partial(\text{local structure})}\approx 0$), forming a more robust optimization target.
TABLE IV: Comparison with Different Loss Functions
Loss Function Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
MSE Loss 14.42 87.00 19.47 70.60
MAE Loss 13.63 85.90 19.29 67.50
Huber Loss 14.64 87.40 19.54 68.40
NPCC Loss 10.36 69.90 17.03 66.90
NSSIM Loss 14.12 85.10 19.41 70.80
Weighted Loss
($\alpha=0.9$)
15.10 88.50 20.04 73.90
Note:The weighted loss combines MSE and NSSIM losses, where $\alpha$ represents the weight of MSE (0.9) and 1-$\alpha$ (0.1) is the weight of NSSIM.
V-C Model Variants
To validate the effectiveness of our physics-informed design, we compared Ptychoformer with two state-of-the-art hybrid CNN-ViT models, adapted for ptychography task. The first is a TransUNet-inspired model [37], which uses a hybrid CNN-Transformer encoder where CNN extracts features and the Transformer processes them further. Its decoder employs a cascading structure with skip connections. The second is a SegFormer-based model [38], which features a hierarchical structure with progressive resolution reduction for multi-scale features. This model replaces the first linear layer in the feed-forward network with a $3\times 3$ convolution. We conducted the same experiment as in Section IV-B2 under noise-free conditions. Table V shows that Ptychoformer outperforms both models. TransUNet and SegFormer’s limitations stem from designs optimized for real-space image processing. The TransUNet variant, while benefiting from the combination of CNN and Transformer, still relies on local feature extraction in its initial stages. This approach is suboptimal for capturing the global coherence information present in diffraction patterns. Similarly, the SegFormer variant’s hierarchical structure, while effective for multi-scale feature extraction in natural images, may lose critical high-frequency information in the context of diffraction patterns due to its progressive downsampling.
TABLE V: Performance comparison of Model Variants.
Method Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
TransUNet[37] 13.57 81.20 18.77 67.80
SegFormer[38] 13.76 83.30 19.12 68.10
Ptychoformer 14.42 87.00 19.41 70.60
VI Conclusion
In this study, we introduced Ptychoformer, a deep learning framework designed for ptychographic retrieval from frequency-domain data. By incorporating physical insights into the model architecture, particularly through the PCA mechanism and dual-branch structure, Ptychoformer demonstrated superior performance over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency. The principles underlying Ptychoformer show promise for extension to other frequency-domain inverse imaging problems, potentially advancing high-resolution retrieval techniques across various scientific disciplines.
Acknowledgments
We thank the University of Sydney’s Digital Sciences Initiative for financial support of this project through grant DSI Ignite grant scheme.

### v2

@@
-Ptychoformer: A Physics-Guided Deep Learning Framework for Ptychographic Imaging
+A Physics-Inspired Deep Learning Framework with Polar Coordinate Attention for Ptychographic Imaging
 Abstract
-Ptychographic imaging confront limitations in applying deep learning (DL) for retrieval from diffraction patterns.
-Conventional neural architectures are optimized for natural images, overlooking the unique physical characteristics of diffraction data, including radial intensity decay and coherent information distributed in concentric rings.
-In this paper, we present Ptychoformer, a physics-guided DL framework for ptychographic imaging that aligns attention mechanisms and feature extraction with these diffraction physics properties through introducing a dual-branch architecture which accounts for both local and non-local dependencies from the patterns.
-It consists of a Polar Coordinate Attention (PCA) mechanism that is inspired by the Ewald construction in X-ray crystallography to enhance high-frequency component fidelity.
-Experimental results demonstrate Ptychoformer’s superior performance across both simulated and real data in preserving fine details and suppressing artifacts.
-On simulated data, Ptychoformer achieves up to 5.4% higher PSNR and 4.2% higher SSIM for amplitude retrieval compared to existing methods.
-For real experimental data, it demonstrates up to 12.5% higher PSNR and 31.3% higher SSIM for amplitude retrieval.
-Notably, Ptychoformer maintains robust performance under limited training data and low overlap ratios, outperforming existing models.
+Ptychographic imaging confronts inherent challenges in applying deep learning for phase retrieval from diffraction patterns.
+Conventional neural architectures, both convolutional neural networks and Transformer-based methods, are optimized for natural images with Euclidean spatial neighborhood-based inductive biases that exhibit geometric mismatch with the concentric coherent patterns characteristic of diffraction data in reciprocal space.
+In this paper, we present PPN, a physics-inspired deep learning network with Polar Coordinate Attention (PoCA) for ptychographic imaging, that aligns neural inductive biases with diffraction physics through a dual-branch architecture separating local feature extraction from non-local coherence modeling.
+It consists of a PoCA mechanism that replaces Euclidean spatial priors with physically consistent radial-angular correlations.
+PPN outperforms existing end-to-end models, with spectral and spatial analysis confirming its greater preservation of high-frequency details.
+Notably, PPN maintains robust performance compared to iterative methods even at low overlap ratios — well-suited for high-throughput imaging in real-world acquisition scenarios for samples with consistent structural characteristics.
 Index Terms:
-Ptychography, Physics-Guided Deep Learning, Inverse Imaging Problems, Transformer.
+Ptychography, Physics-Inspired Deep Learning, Reciprocal-Space Learning, Transformer.
 I Introduction
@@
 Addressing the inherent limitations of existing methods requires a fundamental reconsideration of deep learning architectures for ptychographic retrieval, with a focus on optimizing neural networks for representing diffraction patterns in reciprocal space.
-This insight guided the development of Ptychoformer, a novel framework that bridges frequency-domain diffraction patterns and real-space images using physics-informed deep learning techniques.
+This insight inspired the development of PPN, a novel framework that bridges frequency-domain diffraction patterns and real-space images using physics-informed deep learning techniques.
 Drawing inspiration from the Ewald construction in X-ray crystallography, as depicted in Fig.
@@
 This insight revealed that diffraction patterns are inherently defined in spherical rather than planar geometry.
-Consequently, we developed Polar Coordinate Attention (PCA), an attention mechanism leveraging polar coordinates that naturally aligns with diffraction physics.
+Consequently, we developed Polar Coordinate Attention (PoCA), an attention mechanism leveraging polar coordinates that naturally aligns with diffraction physics.
 This polar coordinate-based attention mechanism reframes real-space structural priors in the diffraction context, mapping scattering vector magnitude to $r$ and angular information to $\theta$, thus capturing both radial intensity attenuation and angular coherence.
 Recognizing the multi-scale and multi-frequency nature of diffraction patterns, as shown in the right side of Fig.
-1, Ptychoformer employs a dual-branch architecture combining a Local Dependencies Branch constructed from standard ViT blocks with a NonLocal Coherence block containing our designed PCA.
+1, PPN employs a dual-branch architecture combining a Local Dependencies Branch constructed from standard ViT blocks with a NonLocal Coherence block containing our designed PoCA.
 This design balances the capture of long-range and local dependencies in diffraction patterns.
-Ptychoformer’s architecture intrinsically aligns with ptychography physics, creating a more natural correspondence between model operations and underlying physical processes.
+PPN’s architecture intrinsically aligns with ptychography physics, creating a more natural correspondence between model operations and underlying physical processes.
 This structural redesign based on physical insights enables our model to more effectively extract and utilize global coherent information and long-range dependencies in diffraction data, thus offering a physics-informed approach to ptychographic retrieval.
-Our comprehensive experiments on both simulated and real datasets demonstrate Ptychoformer’s significant advantages over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency.
-Under challenging conditions including low overlap ratios, and limited training data, our model maintains superior retrieval quality, achieving up to 36.1% reduction in Mean Squared Error (MSE) and 12.03% increase in Structural Similarity Index Measure (SSIM) compared to the best-performing existing end-to-end methods.
+Our comprehensive evaluation across simulated and real experimental datasets demonstrates PPN’s multi-faceted advantages: (1) Superior retrieval quality across all metrics against leading end-to-end baselines, particularly in high-frequency preservation as evidenced through power spectral density curves and cross-sectional intensity profiles; (2) Remarkable operational viability - maintaining $<$5% performance degradation at 30% overlap ratio while achieving $>$1,000× faster inference than iterative method when tested on samples without feature distribution shifts, crucial for time-sensitive synchrotron experiments; (3) Unprecedented efficiency with 11× fewer parameters than transformer-based counterparts (6.1M vs 68.9M) and Floating Point Operations Per Second (FLOPs) comparable to lightweight CNNs, enabling deployment in real-world acquisition scenarios.
 These improvements could significantly enhance the applicability of ptychography in time-sensitive or radiation-sensitive imaging scenarios across various scientific disciplines.
 The primary contributions of this work are summarized as follows:
-  • We present Ptychoformer, an end-to-end deep learning framework addressing limitations of existing DL-based ptychography methods.
-Unlike conventional neural architectures optimized for spatial-domain natural images, our approach incorporates designs aligned with diffraction physics.
-  • We propose the PCA mechanism, a physics-inspired approach that decomposes attention into radial and angular components, enabling efficient extraction of coherent information across frequency bands in diffraction patterns.
-  • Extensive experiments on both simulated and real datasets show that the proposed method achieves high-fidelity and superior artifact reduction, while maintaining efficiency under challenging conditions including noisy environments and limited data acquisition scenarios.
+  • We present PPN, a physics-inspired dual-branch framework specifically designed for ptychographic imaging that addresses the geometric mismatch between Euclidean spatial priors and concentric coherent patterns in reciprocal space.
+  • We propose the PoCA mechanism that replaces spatial neighborhood priors with radial-angular correlations, achieving superior high-frequency preservation compared to other end-to-end baselines.
+  • We demonstrate PPN’s practical advantages with $>$1000× faster inference than iterative method, $<$5% performance degradation at 30% overlap ratio when tested on samples without feature distribution shifts.
 II Related Works
@@
 The physics-informed automatic differentiation ptychography (ADP) framework uses pre-trained autoencoders to map high-dimensional image data to a low-dimensional latent space [15], while the double deep image prior (DDIP) method reduces the optimization parameter space [16], both enhancing convergence rates and noise robustness.
+However, these methods still require traditional iterative algorithms.
 II-A2 Post-processing
-Neural networks refine reconstructions from traditional algorithms, including enhancing a single iteration of the Gerchberg-Saxton-Fienup (GSF) algorithm to improve spatial resolution and reduce artifacts [5], showing promise in live cell imaging.
-II-A3 End-to-End
+Neural networks refine reconstructions from traditional algorithms, including enhancing a single iteration of the iterative algorithm to improve spatial resolution and reduce artifacts [5] [17].
+While these methods significantly improve the quality of reconstructed images, they still rely on initial reconstructions and cannot provide real-time or fully automated solutions.
+II-A3 End-to-End with CNNs
 End-to-end methods directly map diffraction patterns to complex object functions, bypassing iterative processes.
 Examples include PtychoNN with a modified U-Net and two-branch decoder for amplitude and phase [9], PtyNet with group convolution and Leaky ReLU for efficiency [11], and DPI using a traditional U-Net with skip connections [10].
-Despite improved retrieval speed, these methods face challenges to preserve high-frequency information and the global coherence of diffraction patterns, as CNNs and standard Vision Transformers (ViTs) are tailored for real-space images with local correlations.
-Recent studies using ViTs for ptychographic retrieval leverage self-attention for long-range dependencies [17], but applying ViT architectures to frequency-domain data is limited by radial intensity decay and concentric information distribution.
-Addressing these frequency-domain challenges, including processing global coherence and preserving high frequencies, remains crucial for advancing retrieval quality and efficiency, especially with limited data or low overlap ratios.
-II-B ViT-based Methods
-Figure 2: The proposed Ptychoformer architecture for ptychographic imaging.
+Whereas CNNs excel at local feature extraction for natural images[18, 19], their inductive biases prove suboptimal for diffraction patterns requiring global phase coherence.
+The quadratic phase factors in Fresnel propagation create position-dependent correlations that span the entire detector plane, yet the local receptive fields of CNNs (typically 3×3) cannot span the full radial extent of diffraction rings, causing these physically critical phase relationships between distant pixels to be irrecoverably lost during feature encoding.
+Furthermore, progressive downsampling in hierarchical architectures systematically discards high-frequency information during feature abstraction.
+These inherent limitations of CNNs motivate the exploration of attention mechanisms for better long-range dependency modeling.
+II-B Transformer-based Methods
+Figure 2: The proposed PPN for ptychographic imaging.
 It features dual branches: a Local Dependencies Branch with standard ViT blocks, and a NonLocal Coherence Branch with Polar Coordinate Attention mechanism.
-The model processes logarithmically mapped diffraction patterns, combining features from both branches before decoding into amplitude and phase images with physical constraints applied.
+The model processes logarithmically mapped diffraction patterns, combining features from both branches before decoding into individual amplitude and phase reconstructions at each position, which are then separately stitched to generate full field of view images for both amplitude and phase.
+Vision Transformers (ViTs) [20] emerged as a promising solution to the limitations of CNNs, demonstrating exceptional capability in capturing long-range dependencies across various domains [21, 22, 23, 24, 25].
+Recent adaptations in ptychography, such as PtychoFormer[26] employs a hybrid architecture combining ViT and CNN components, specifically a SegFormer variant adapted for ptychography.
+While it surpasses traditional CNN methods in reconstruction accuracy, this comes with significant increases in parameter count and training costs.
+PtychoDV [17] adopts a compromise strategy that utilizes ViTs’ reconstruction results as initial guesses for iterative methods, thereby enhancing the final accuracy of conventional iterative approaches.
+This indirectly reveals that pure end-to-end deep learning methods still face performance bottlenecks in diffraction reconstruction tasks - the core issue we aim to address in this paper.
+While Transformers demonstrate superior capability in capturing global contexts, their direct application to diffraction patterns neglects crucial physical priors inherent in reciprocal space representations.
 Diffraction patterns, governed by wave optics, approximate the Fourier transform of an object’s transmission function in far-field conditions [27], resulting in reciprocal space features like sparse representations and concentric rings [28].
 Accurate image retrieval and phase retrieval require effectively capturing both sparsity and long-range dependencies within these patterns.
-Traditional CNNs, though effective for natural images [20, 21], are limited with the global nature of diffraction phenomena due to their local receptive fields.
-Vision Transformers (ViTs), introduced by Dosovitskiy et al.
-[22], offer a promising alternative with self-attention mechanisms that capture global features and long-range dependencies [23], outperforming CNNs in various computer vision tasks [24, 25] and extending to fields like medical image segmentation [24, 26], climate modeling [27], and molecular structure prediction [28].
-However, applying ViTs to phase retrieval in CDI is challenging due to the lack of CNN-like inductive biases such as translational invariance and local connectivity, coupled with often insufficient data in practical phase imaging.
-For example, PtychoDV [17] uses ViTs for feature extraction from diffraction patterns but shows inferior retrieval performance compared to CNN-based U-Net architectures without iterative post-processing.
-This gap highlights the potential for developing hybrid methods that leverage transformers’ global feature extraction while addressing data limitations in ptychography.
+Although Vision Transformers achieve global modeling through multi-head self-attention mechanisms (MHSA), their Euclidean spatial neighborhood-based attention weight calculation exhibits geometric mismatch with the concentric coherent patterns in reciprocal space characteristic of diffraction patterns.
+Our focus is to resolve the critical bottleneck of effective information extraction from diffraction patterns and improved reconstruction of high-frequency spatial information, rather than surpassing or replacing the ultimate resolution achieved by iterative algorithms.
 III Methodology
@@
 To account for real-world factors, we model the diffraction intensity as:
-$\begin{split}I_{j}(\mathbf{q})=&\mathcal{P}_{\eta}\left(\eta\left|\mathcal{F}\left\{\mu P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})O(\mathbf{r})\right.\right.\right.\\ &\left.\left.\left.+(1-\mu)P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})\mathbb{E}_{\mathbf{r}}[O(\mathbf{r})]\right\}\right|^{2}\right)\\ &+\mathcal{N}(0,\sigma^{2})\end{split}$ (2)
-where $\mu$ is the coherence parameter, $\eta$ is the detector efficiency, $\mathcal{P}_{\eta}$ models Poisson shot noise, $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise, and $\delta\mathbf{r}_{j}$ accounts for positional jitter.
-The objective function for ptychographic retrieval is defined as:
+$\begin{split}I_{j}(\mathbf{q})=&\text{Poisson}\Big{(}\eta\Big{|}\mathcal{F}\Big{\{}\mu P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})O(\mathbf{r})\\ &+(1-\mu)P(\mathbf{r}-\mathbf{r}_{j}-\delta\mathbf{r}_{j})\cdot\mathbb{E}_{\mathbf{r}}[O(\mathbf{r})]\Big{\}}\Big{|}^{2}\Big{)}\\ &+\mathcal{N}(0,\sigma^{2})\end{split}$ (2)
+where $\mu$ is the coherence parameter, $\eta\in[0,1]$ is the detector efficiency representing the quantum efficiency (probability of detecting each photoelectron), $\text{Poisson}(\lambda)$ denotes the Poisson distribution with rate parameter $\lambda=\eta\cdot\left|\mathcal{F}\left\{\dots\right\}\right|^{2}$, where |{}$\left|\mathcal{F}\left\{\dots\right\}\right|^{2}$ is the theoretical diffraction intensity (in photons/pixel) and their product forms the expected value of the Poisson process, and $\delta\mathbf{r}_{j}$ accounts for positional jitter.
+Based on the physical model of the actual imaging process in Eq. (2), we formulate the objective function for ptychographic retrieval in Eq. (3) to effectively recover the object and probe functions:
 $\begin{split}\mathcal{J}(O,P)=&\frac{1}{J}\sum_{j=1}^{J}\left\|I_{j}-|\mathcal{F}\{P(\mathbf{r}-\mathbf{r}_{j})\cdot O(\mathbf{r})\}|^{2}\right\|_{2}^{2}\\ &+\lambda_{1}\Omega_{1}(O)+\lambda_{2}\Omega_{2}(P)\end{split}$ (3)
-where $\left\|\mathbf{x}\right\|_{2}^{2}=\sum_{i}x_{i}^{2}$ is the squared Euclidean norm, $\Omega_{1}(O)=\left\|\nabla O\right\|_{1}$ enforces sparsity in the object gradient, $\Omega_{2}(P)=\left\|P-P_{0}\right\|_{2}^{2}$ constrains the probe function to maintain proximity to to an initial estimate $P_{0}$, and $\lambda_{1}$ and $\lambda_{2}$ are regularization weights.
-III-B Ptychoformer Architecture
-Figure 3: Architectural overview of the PCA mechanism.
-The diagram illustrates the flow of input features through polar coordinate transformation, query-key-value attention computations, and subsequent operations including polar coordinate weighting and Hadamard product.
-The process culminates in multi-head attention, followed by concatenation and linear projection to produce the final output.
-The Ptychoformer architecture uses a bifurcated structure to map diffraction patterns to a reconstructed complex-valued object function.
+where $\left\|\mathbf{x}\right\|_{2}^{2}=\sum_{i}x_{i}^{2}$ is the squared Euclidean norm, $\Omega_{1}(O)=\left\|\nabla O\right\|_{1}$ enforces sparsity in the object gradient, $\Omega_{2}(P)=\|P-P_{0}\|_{F}^{2}$ constrains the probe function to maintain proximity to to an initial estimate $P_{0}$, and $\lambda_{1}$ and $\lambda_{2}$ are regularization weights.
+III-B PPN Architecture
+The proposed architecture employs a bifurcated structure to map diffraction patterns to a reconstructed complex-valued object function.
 The input undergoes a logarithmic transformation $I_{\text{log}}(\mathbf{q},j)=\log(1+I_{j}(\mathbf{q}))$ as a preprocessing step to address the large disparity between high and low frequency values in the diffraction patterns.
 III-B1 Local Dependencies Branch
-The Local Dependencies Branch employs a standard ViT to analyze diffraction patterns, aiming to capture local features and relationships within the pattern.
-The branch’s structure consists of several key components.
-Initially, a patch extraction mechanism segments the diffraction pattern into smaller sections.
-These segments then undergo reshaping to facilitate subsequent processing.
-Following this, an encoding step is applied to each patch, designed to emphasize relevant features.
-The encoded patches are then processed through a series of ViT blocks.
-Each ViT block $\mathcal{B}_{l}$ in $\mathcal{B}_{L}$ is defined as:
-$\begin{split}\mathcal{B}_{l}(\mathbf{X})&=\mathbf{X}+\mathcal{FFN}(\mathcal{LN}(\mathbf{X}+\mathcal{MSA}(\mathcal{LN}(\mathbf{X}))))\\ \mathcal{MSA}(\mathbf{X})&=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})\mathcal{W}^{O}\\ \text{head}_{i}&=\text{Attention}(\mathbf{X}\mathcal{W}^{Q}_{i},\mathbf{X}\mathcal{W}^{K}_{i},\mathbf{X}\mathcal{W}^{V}_{i})\end{split}$ (4)
-where $\mathcal{LN}$ denotes layer normalization, $\mathcal{MSA}$ is multi-head self-attention with $h$ heads, $\mathcal{FFN}$ is a position-wise feed-forward network, and $\mathcal{W}^{Q}_{i},\mathcal{W}^{K}_{i},\mathcal{W}^{V}_{i},\mathcal{W}^{O}$ are learnable weight matrices.
-Despite apparent disruption, patch extraction facilitates low-frequency information capture, especially from central patches containing crucial low-frequency components.
-III-B2 NonLocal Coherence Branch
-The NonLocal Coherence Branch operates directly on the full-resolution input, intentionally omitting patch extraction to preserve the global structure of the diffraction pattern.
-It replaces the multi-head self-attention in standard ViT blocks with PCA, and the radial information introduced by PCA eliminates the need for position embeddings.
-While traditional self-attention mechanisms are effective for capturing long-range dependencies, they are not optimized for ptychographic data.
-Designed for natural images, they incorporate spatial positional encodings $\text{PE}(p)$ that can bias attention towards local regions: $A_{pq}=\text{softmax}(\frac{(x_{p}+\text{PE}(p))^{T}(x_{q}+\text{PE}(q))}{\sqrt{d_{k}}})$.
-This provides structural priors suited for conventional visual tasks but may not be optimal for frequency-domain diffraction patterns.
-To address this limitation, we introduce the PCA mechanism.
-PCA transforms spatial coordinates into polar coordinates, shifting the structural priors from Cartesian spatial locality to the concentric information distribution inherent in frequency-domain diffraction patterns.
-This transformation enables PCA to better capture the radial and angular dependencies characteristic of ptychographic data.
-The PCA layer $\mathcal{PCA}:\mathbb{R}^{B\times H\times W\times C}\to\mathbb{R}^{B\times H\times W\times C^{\prime}}$ is defined as:
-$\mathcal{PCA}(\mathbf{X})=\text{Reshape}(\mathcal{F}(\text{Flatten}(\mathbf{X})))$ (5)
-$\mathcal{F}(\mathbf{X})=\text{Concat}(\text{Head}_{1},\text{Head}_{2},\dots,\text{Head}_{h})\mathcal{W}_{O}$ (6)
-$\text{Head}_{i}=\text{softmax}(\mathcal{A}r^{i}(\mathbf{X})+\mathcal{A}\theta^{i}(\mathbf{X}))\mathcal{V}^{i}(\mathbf{X})$ (7)
-$\mathcal{A}r^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{R}$ (8)
-$\mathcal{A}\theta^{i}(\mathbf{X})=\frac{\mathcal{Q}^{i}(\mathbf{X})\mathcal{K}^{i}(\mathbf{X})^{T}}{\sqrt{d_{k}}}\odot\mathbf{\Theta}$ (9)
-where $\mathcal{Q}^{i},\mathcal{K}^{i}\in\mathbb{R}^{N\times d_{k}}$, $\mathcal{V}^{i}\in\mathbb{R}^{N\times d_{v}}$ are the linear projections for queries, keys, and values for the $i$-th head, $d_{k}$ is the dimension of the keys, $\mathbf{R},\mathbf{\Theta}\in\mathbb{R}^{N\times N}$ are the normalized radial and angular coordinate matrices, respectively, $\odot$ denotes the Hadamard product (element-wise multiplication), and $\mathcal{W}_{O}\in\mathbb{R}^{hd_{v}\times C^{\prime}}$ is the output projection matrix.
-Here, $N=H\times W$ is the number of spatial positions.
-The coordinate matrices are computed as:
-$R_{ij}=\frac{\sqrt{(i-W/2)^{2}+(j-H/2)^{2}}}{\sqrt{(W/2)^{2}+(H/2)^{2}}}$ (10)
-$\Theta_{ij}=\frac{\arctan 2(j-H/2,i-W/2)}{\pi}$ (11)
-where $i\in[1,W]$ and $j\in[1,H]$ are the horizontal and vertical indices of the spatial positions, respectively.
-This mechanism decomposes attention into radial and angular components, mirroring the symmetry of diffraction patterns.
-Operating without patch extraction, this branch preserves global structure.
-Its polar coordinate attention is designed to extract global coherence information, particularly high-frequency components requiring long-range dependency modeling.
+The Local Dependencies Branch utilizes a standard Vision Transformer (ViT) to analyze diffraction patterns, capturing local features and structural relationships within each pattern.
+Operating on input tensors $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ (where $B$ is batch size, $H\times W$ is spatial resolution, and $C$ is channel depth), this branch processes data through several key components:
+Step 1.
+Patch Extraction divides the input image into non-overlapping patches of size $P\times P$, resulting in $\mathbf{X}_{\text{patches}}\in\mathbb{R}^{B\times N_{p}\times(P^{2}\cdot C)}$ where $N_{p}=\frac{H}{P}\times\frac{W}{P}$ represents the total number of patches.
+Step 2.
+Linear Projection maps each patch to a $D$-dimensional embedding space using $\mathbf{X}_{\text{embed}}=\mathbf{X}_{\text{patches}}\mathbf{W}_{E}\in\mathbb{R}^{B\times N_{p}\times D}$ where $\mathbf{W}_{E}\in\mathbb{R}^{(P^{2}C)\times D}$ is the projection matrix.
+Step 3.
+Positional Encoding adds spatial information via $\mathbf{X}_{\text{pos}}=\mathbf{X}_{\text{embed}}+\mathbf{E}_{\text{pos}}$, where the positional encoding matrix $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{N_{p}\times D}$ is broadcasted across the batch dimension to match $\mathbf{X}_{\text{embed}}\in\mathbb{R}^{B\times N_{p}\times D}$.
+Step 4.
+Transformer Blocks with Pre-LN structure process the sequence through $L$ layers according to $\mathbf{X}^{\prime(l)}=\mathcal{MSA}(\mathcal{LN}(\mathbf{X}^{(l-1)}))+\mathbf{X}^{(l-1)}$ and $\mathbf{X}^{(l)}=\mathcal{FFN}(\mathcal{LN}(\mathbf{X}^{\prime(l)}))+\mathbf{X}^{\prime(l)}$.
+The Multi-Head Self-Attention (MSA) mechanism computes $\mathcal{MSA}(\mathbf{X})=\text{Concat}(\text{head}_{1},...,\text{head}_{h})\mathbf{W}^{O}$ where each attention head captures different relationship patterns in the data, with projection matrices $\mathbf{W}^{Q}_{i},\mathbf{W}^{K}_{i}\in\mathbb{R}^{D\times d_{k}}$ and $\mathbf{W}^{V}_{i}\in\mathbb{R}^{D\times d_{v}}$ (typically $d_{k}=d_{v}=D/h$).
+Step 5. Spatial Restoration reshapes the output to $\mathcal{H}_{\text{Local}}\in\mathbb{R}^{B\times(H/P)\times(W/P)\times D}$.
+[…]
+III-B2 Non-Local Coherence Branch
+The Non-Local Coherence Branch employs an architectural framework similar to the Local Dependencies Branch, sharing identical transformer block components while introducing our novel Polar Coordinate Attention (PoCA) mechanism.
+This specialized attention mechanism processes input at pixel-level granularity ($H\times W$ pixels) without patch extraction, enabling the capture of long-range dependencies across diffraction patterns.
+The PoCA mechanism systematically encodes physical constraints inherent to diffraction physics, including the $I\propto r^{-4}$ intensity distribution through radial decay weighting, preserves crystallographic symmetry via angular continuity, and adapts to illumination shifts with dynamically learned centering parameters $\alpha_{x}$ and $\alpha_{y}$.
+After logarithmic transformation, the single-channel diffraction pattern $\mathbb{R}^{B\times H\times W\times 1}$ is projected to a multi-channel tensor $\mathbb{R}^{B\times H\times W\times C}$ via a learnable $1\times 1$ convolutional layer, where $C$ denotes the embedding dimension.
+For an input tensor $\mathbf{X}\in\mathbb{R}^{B\times H\times W\times C}$ (where $B$ represents batch size, $H\times W$ spatial resolution), this branch processes data through several key components:
+Step 1.
+Tensor Flattening transforms the input into a sequence representation via $\mathbf{X}_{f}=\text{Reshape}(\mathbf{X},(B,H\times W,C))\in\mathbb{R}^{B\times N\times C}$, where $N=H\times W$ represents the total number of pixels.
+Step 2. Query-Key-Value Projection maps the flattened tensor into three separate embedding spaces using:
+$\mathcal{Q}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{Q}^{i}\quad$ $\mathbf{W}_{Q}^{i}$ $\in\mathbb{R}^{C\times d_{k}}$ (4)
+$\mathcal{K}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{K}^{i}\quad$ $\mathbf{W}_{K}^{i}$ $\in\mathbb{R}^{C\times d_{k}}$
+$\mathcal{V}^{i}$ $=\mathbf{X}_{f}\mathbf{W}_{V}^{i}\quad$ $\mathbf{W}_{V}^{i}$ $\in\mathbb{R}^{C\times d_{v}}$
+where projection matrices $\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i}\in\mathbb{R}^{C\times d_{k}}$ and $\mathbf{W}_{V}^{i}\in\mathbb{R}^{C\times d_{v}}$ define the embedding transformations for attention head $i$.
+The resulting tensors are $\mathcal{Q}^{i},\mathcal{K}^{i}\in\mathbb{R}^{B\times N\times d_{k}}$ and $\mathcal{V}^{i}\in\mathbb{R}^{B\times N\times d_{v}}$, where the matrix multiplication is performed along the feature dimension $C$.
+Polar coordinate parameterization defines the physical diffraction geometry using:
+$c_{x}=\frac{W}{2}+\alpha_{x}\frac{W}{2},\quad c_{y}=\frac{H}{2}+\alpha_{y}\frac{H}{2}$ (5)
+$r_{m}=\frac{\log(1+\|\mathbf{p}_{m}-\mathbf{c}\|)}{\log(1+r_{\max})},\quad\theta_{m}=\arctan 2\left(y_{m}-c_{y},x_{m}-c_{x}\right)$ (6)
+$\mathbf{\Phi}_{r}^{mn}=\frac{1}{1+|r_{m}-r_{n}|},\quad\mathbf{\Phi}_{\theta}^{mn}=\cos(\theta_{m}-\theta_{n})$ (7)
+where $r_{\max}=\sqrt{(W/2)^{2}+(H/2)^{2}}$ ensures normalized radial coordinates, $\mathbf{p}_{m}=(x_{m},y_{m})$ denotes original pixel coordinates, $\mathbf{c}=(c_{x},c_{y})$ represents the learned diffraction center, and $\alpha_{x},\alpha_{y}\in[-0.5,0.5]$ are learnable parameters initialized at 0.
+The implementation employs logarithmic scaling to compress the dynamic range of radial distances, with angle normalization $\theta_{m}\in[0,2\pi)$ ensuring continuity.
+Step 3. Polar Coordinate Modulation incorporates physical constraints by modulating the standard dot-product attention with radial and angular weighting matrices:
+$\begin{split}\mathcal{A}_{\text{polar}}^{i}&=\underbrace{\frac{\mathcal{Q}^{i}(\mathcal{K}^{i})^{\top}}{\sqrt{d_{k}}}}_{\text{base attention}}\odot\underbrace{\mathbf{\Phi}_{r}}_{\text{radial decay}}\odot\underbrace{\mathbf{\Phi}_{\theta}}_{\text{angular continuity}}\end{split}$ (8)
+where $\mathbf{\Phi}_{r}$, $\mathbf{\Phi}_{\theta}\in\mathbb{R}^{1\times N\times N}\quad\text{(broadcastable to }\mathbb{R}^{B\times N\times N}\text{)}$, $\odot$ represents element-wise multiplication with broadcasting, imposing both radial decay and angular continuity constraints.
+Step 4.
+Attention Application computes each attention head through $\text{Head}_{i}=\text{softmax}\left(\mathcal{A}_{\text{polar}}^{i}\right)\mathcal{V}^{i}$, where the softmax normalization ensures proper probability distribution across the attention weights, and $\mathcal{A}_{\text{polar}}^{i}\in\mathbb{R}^{B\times N\times N}$ is applied to $\mathcal{V}^{i}\in\mathbb{R}^{B\times N\times d_{v}}$.
+Step 5.
+Multi-head Integration combines information from all h attention heads via $\mathcal{G}(\mathbf{X}_{f})=\text{Concat}(\text{Head}_{1},...,\text{Head}_{h})\mathbf{W}^{O}$, with $\mathbf{W}^{O}\in\mathbb{R}^{h\cdot d_{v}\times C^{\prime}}$ serving as the output projection matrix and $\text{Head}_{i}\in\mathbb{R}^{B\times N\times d_{v}}$.
+Step 6.
+Spatial Restoration reshapes the processed output to match the original spatial dimensions through $\mathcal{P}\text{o}CA{}(\mathbf{X})=\text{Reshape}\left(\mathcal{G}(\mathbf{X}_{f}),(B,H,W,C^{\prime})\right)\in\mathbb{R}^{B\times H\times W\times C^{\prime}}$.
+[…]
+By decomposing attention into radial and angular components, PoCA effectively mimics the Ewald sphere’s intersection with the reciprocal lattice (Fig.
+1(b)), where $r_{mn}$ correlates with the reciprocal space resolution of scattering vectors, and $\theta_{mn}$ encodes Bragg angle relationships through angular coherence constraints.
 III-B3 Feature Fusion and Decoding
-Let $\mathcal{F}_{\text{Local}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}:L^{2}(\mathbb{R}^{2}\times\{1,\ldots,J\})\to\mathcal{H}_{\text{NonLocal}}$ denote the feature extraction operators from the Local Dependencies Branch and the NonLocal Coherence Branch, respectively.
-The decoder is defined as a composition of three upsampling operators $\mathcal{D}=\mathcal{D}_{3}\circ\mathcal{D}_{2}\circ\mathcal{D}_{1}$, each progressively increasing the spatial resolution using transposed convolutions.
-Here, $\circ$ denotes function composition, such that $(f\circ g)(x)=f(g(x))$ for any compatible functions $f$ and $g$.
-The feature fusion process, combining features from both branches, is formulated as:
-$\mathcal{F}_{\text{fused}}=\mathcal{C}(\mathcal{F}_{\text{Local}},\mathcal{F}_{\text{NonLocal}})$ (12)
-where $\mathcal{C}$ concatenates $\mathcal{F}_{\text{Local}}$ and $\mathcal{F}_{\text{NonLocal}}$, followed by a convolution to integrate the features.
-The fused features $\mathcal{F}_{\text{fused}}$ are then processed through the three-layer decoder $\mathcal{D}$, implemented using transposed convolutions with batch normalization and ReLU activations, split into parallel branches $\mathcal{D}_{\text{amplitude}}$ and $\mathcal{D}_{\text{phase}}$, to reconstruct the preliminary amplitude $\mathbf{A}(\mathbf{r})$ and phase $\boldsymbol{\phi}(\mathbf{r})$ images.
-III-B4 Output Constraints
-To ensure the physical validity of the reconstructed complex-valued object function, we apply constraints to the final outputs.
-The amplitude $\mathbf{A}_{\text{final}}:\mathbb{R}^{2}\to\mathbb{R}^{+}$ and phase $\boldsymbol{\phi}_{\text{final}}:\mathbb{R}^{2}\to[-\pi,\pi]$ are constrained as:
-$\mathbf{A}_{\text{final}}(\mathbf{r})=A_{\text{max}}\cdot\sigma\left(\frac{\mathbf{A}(\mathbf{r})}{A_{\text{max}}}\right),\quad\boldsymbol{\phi}_{\text{final}}(\mathbf{r})=\pi\cdot\tanh\left(\boldsymbol{\phi}(\mathbf{r})\right)$ (13)
-where $A_{\text{max}}$ is the maximum allowable amplitude, $\sigma$ is the sigmoid function, and $\tanh$ is the hyperbolic tangent function.
-The final complex-valued object function is reconstructed as
-$O_{\text{reconstructed}}(\mathbf{r})=\mathbf{A}_{\text{final}}(\mathbf{r})\cdot\exp\left(i\boldsymbol{\phi}_{\text{final}}(\mathbf{r})\right).$
-III-B5 Training Objective
+The feature fusion process addresses the resolution mismatch between the two branches.
+The Local Dependencies Branch output $\mathcal{H}_{\text{Local}}\in\mathbb{R}^{B\times(H/P)\times(W/P)\times D}$ has lower spatial resolution than the Non-Local Branch output $\mathcal{H}_{\text{NonLocal}}\in\mathbb{R}^{B\times H\times W\times C^{\prime}}$.
+To align these representations, we employ bilinear upsampling on the Local Branch features:
+$\mathcal{H}_{\text{Local}}^{\text{upsampled}}=\text{Upsample}_{\text{bilinear}}(\mathcal{H}_{\text{Local}})\in\mathbb{R}^{B\times H\times W\times D}$ (9)
+The aligned features are then combined through concatenation followed by a 1×1 convolution:
+$\mathcal{H}_{\text{fused}}=\text{Conv}_{1\times 1}(\text{Concat}[\mathcal{H}_{\text{Local}}^{\text{upsampled}},\mathcal{H}_{\text{NonLocal}}])$ (10)
+The decoder $\mathcal{D}$ consists of three transposed convolution blocks with progressive upsampling:
+$\begin{split}\mathcal{D}_{i}&=\text{BatchNorm}(\text{ReLU}(\text{ConvTranspose2D}_{k,s}))\\ \mathcal{D}&=\mathcal{D}_{3}\circ\mathcal{D}_{2}\circ\mathcal{D}_{1}\end{split}$ (11)
+where $k$ is kernel size and $s$ is stride.
+The decoder branches into parallel paths $\mathcal{D}_{\text{amplitude}}$ and $\mathcal{D}_{\text{phase}}$ to reconstruct the object’s amplitude and phase components.
+III-B4 Training Objective
 The loss function $\mathcal{L}:\Theta\to\mathbb{R}^{+}$ is defined as:
 $\mathcal{L}(\theta)=\frac{1}{N}\sum_{n=1}^{N}\left[\left\|A_{\text{true}}^{(n)}-A_{\text{pred}}^{(n)}\right\|_{2}^{2}+\left\|\phi_{\text{true}}^{(n)}-\phi_{\text{pred}}^{(n)}\right\|_{2}^{2}\right]$ (12)
-where $N$ is the number of training samples, $A_{\text{true}}^{(n)}$ and $\phi_{\text{true}}^{(n)}$ are the true amplitude and phase for the $n$-th sample, respectively, and $A_{\text{pred}}^{(n)}=\mathbf{A}_{\text{final}}^{(n)}$ and $\phi_{\text{pred}}^{(n)}=\boldsymbol{\phi}_{\text{final}}^{(n)}$ are the corresponding model predictions after applying the output constraints.
-The optimization problem is formulated as:
+where $N$ is the number of training samples, $A_{\text{true}}^{(n)}$ and $\phi_{\text{true}}^{(n)}$ are the true amplitude and phase for the $n$-th sample (both amplitude and phase values are normalized to the range [0,1]), respectively, and $A_{\text{pred}}^{(n)}=A_{\text{final}}^{(n)}$ and $\phi_{\text{pred}}^{(n)}=\boldsymbol{\phi}_{\text{final}}^{(n)}$ are the corresponding model predictions.
+The two error terms naturally maintain comparable scales without requiring additional weighting factors.The optimization problem is formulated as:
 $\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)$ (13)
@@
 IV-A Experiment Setup
+Figure 3: Performance comparison of single-shot experiment results and full-stitched scene retrieval using simulated data.
+(a) and (b) show amplitude and phase reconstructions at two representative scan positions.
+(c) displays the full-field amplitude and phase images stitched together from individual reconstructions at all scan positions across different models.
 IV-A1 Implementation Details
-The model was trained using a batch size of 32 for 25 epochs.
-Adam optimizer with ReduceLROnPlateau callback (factor 0.5, patience 2, $\min_{\operatorname{lr}}$ 0.0001) adjusts learning rate based on validation loss.
-Training on Tesla T4 GPU.
-Custom callback monitors memory, time, and convergence.
-Model compiled with MSE loss.
-ReduceLROnPlateau strategy implemented.
-Dataset shuffled with 5% validation split.
+We implemented both our DL-based method and the traditional iterative algorithm (like the extended ptychographic iterative engine (ePIE) algorithm[29]) in Python with CUDA acceleration on the same hardware platform.
+The iterative algorithm was implemented with CUDA support, initialized with known probe measurements and processing 32$\times$32 diffraction patterns, with phase unwrapping applied to retrieve the continuous phase.
+The ePIE implementation typically required 100-200 iterations to converge and 75-80% spatial overlap, with minimal gains beyond.
+Both implementations were trained and evaluated on a NVIDIA® Tesla T4 GPU with Intel® Xeon® Platinum 8480CL CPU @ 2.00GHz running Linux 6.1.85+.
+For the DL-based method, we used TensorFlow with an MSE loss function and the Adam optimizer.
+We set the initial learning rate to 1.0 and implemented a ReduceLROnPlateau callback (factor 0.5, patience 2, $\min_{\operatorname{lr}}$ 0.0001) to dynamically adjust the learning rate based on validation loss.
+We used a batch size of 32 with 5% validation split, while a custom callback monitored system memory usage, computation time, and convergence behavior.
 IV-A2 Datasets
-We utilize two types of datasets in our experiments: Simulated Dataset: Dataset based on magnetic nanostructure simulation [29] using Ni80Fe20 parameters.
-512×512 LTEM images represent 5×5 nm areas.
-The dataset contains 2982 images with 75% overlap and 3-pixel jitter.
+We utilize two types of datasets in our experiments: Simulated Dataset: The dataset consists of a 512×512 LTEM image[30].
+The pattern is based on a slightly modified version of the 1951 USAF resolution test chart.
+Using a probe size of 32×32 pixels with a 75% overlap rate and 3-pixel jitter, we generated a total of 3721 patches arranged in a 61×61 grid, each patch being 32×32 pixels.
 Gaussian noise simulates realism, with added Poisson and Gaussian noise for readout and photon effects.
@@
 The validation set is created using 5% of the training data.
+Results on simulated data are presented in Sections IV.B, D, E, F, G and Section V.B, while Section IV.C validates the model using real experimental synchrotron data.
 IV-A3 Evaluation Metrics
 The performance of the models is evaluated using MSE, Peak Signal-to-Noise Ratio (PSNR), and SSIM, with the ground truth values of the materials basis images as references.
-Figure 4: Performance comparison of single-shot experiment results and full-stitched scene retrieval.
-(a) and (b) show two representative scan points.
-(c) represents the full-stitched scene retrieval across models.
-TABLE I: Quantitative comparison of full-stitched scenes under noise-free and noisy conditions. Results are presented as mean $\pm$ standard deviation from ten independent experiments.
+TABLE I: Quantitative comparison of full-stitched scenes under noise-free and noisy conditions on simulated data.
+Results are presented as mean $\pm$ standard deviation from ten independent experiments.
 Condition Method Amplitude Phase
 MSE (×$\times 10^{-2}$) $\downarrow$ PSNR (dB) $\uparrow$ SSIM (%) $\uparrow$ MSE (×$\times 10^{-2}$) $\downarrow$ PSNR (dB) $\uparrow$ SSIM (%) $\uparrow$
-Noise-free PtychoNN 4.32 $\pm$ 0.02 13.52 $\pm$ 0.17 83.50 $\pm$ 0.80 1.22 $\pm$ 0.01 19.09 $\pm$ 0.19 67.10 $\pm$ 0.70
-PtyNet 4.46 $\pm$ 0.03 13.56 $\pm$ 0.15 83.10 $\pm$ 0.60 1.28 $\pm$ 0.01 19.10 $\pm$ 0.26 67.80 $\pm$ 1.30
-DPI 4.29 $\pm$ 0.03 13.68 $\pm$ 0.12 83.40 $\pm$ 0.50 1.26 $\pm$ 0.01 18.91 $\pm$ 0.24 67.70 $\pm$ 1.10
-Ptychoformer 3.63 $\pm$ 0.04 14.42 $\pm$ 0.08 87.00 $\pm$ 0.30 1.14 $\pm$ 0.01 19.41 $\pm$ 0.05 70.60 $\pm$ 0.70
+Noise-free PtychoNN[9] 4.32 $\pm$ 0.02 13.52 $\pm$ 0.17 83.50 $\pm$ 0.80 1.22 $\pm$ 0.01 19.09 $\pm$ 0.19 67.10 $\pm$ 0.70
+PtyNet[11] 4.46 $\pm$ 0.03 13.56 $\pm$ 0.15 83.10 $\pm$ 0.60 1.28 $\pm$ 0.01 19.10 $\pm$ 0.26 67.80 $\pm$ 1.30
+DPI[10] 4.29 $\pm$ 0.03 13.68 $\pm$ 0.12 83.40 $\pm$ 0.50 1.26 $\pm$ 0.01 18.91 $\pm$ 0.24 67.70 $\pm$ 1.10
+PPN (Ours) 3.63 $\pm$ 0.04 14.42 $\pm$ 0.08 87.00 $\pm$ 0.30 1.14 $\pm$ 0.01 19.41 $\pm$ 0.05 70.60 $\pm$ 0.70
 Noisy PtychoNN 6.66 $\pm$ 0.03 11.76 $\pm$ 0.12 74.00 $\pm$ 0.89 1.47 $\pm$ 0.04 18.33 $\pm$ 0.17 58.90 $\pm$ 1.00
@@
 DPI 6.46 $\pm$ 0.02 11.76 $\pm$ 0.10 74.20 $\pm$ 0.65 1.48 $\pm$ 0.06 18.31 $\pm$ 0.20 62.00 $\pm$ 1.20
-Ptychoformer 5.91 $\pm$ 0.03 12.34 $\pm$ 0.11 78.80 $\pm$ 0.59 1.31 $\pm$ 0.06 18.76 $\pm$ 0.16 68.30 $\pm$ 1.10
+PPN (Ours) 5.91 $\pm$ 0.03 12.34 $\pm$ 0.11 78.80 $\pm$ 0.59 1.31 $\pm$ 0.06 18.76 $\pm$ 0.16 68.30 $\pm$ 1.10
 Note: For noisy conditions, noise is simulated using Gaussian and Poisson distributions to model readout noise and photon noise, respectively.
@@
 IV-B Performance Comparison and Analysis
-We evaluated Ptychoformer against three mainstreamed CNN-based methods (PtychoNN, PtyNet, and DPI).
+We evaluated PPN against three mainstreamed CNN-based methods (PtychoNN, PtyNet, and DPI).
 Our analysis covered single-shot retrieval and full-stitched field retrieval under both ideal and noisy conditions.
 IV-B1 Single Scan Point Retrieval Analysis
-Ptychoformer produced retrievals with improved edge definition and clarity compared to those from the other methods, particularly evident in the vertical edge on the left side of Fig.
-4(a).
+PPN reconstructed retrievals with improved edge definition and clarity compared to those from the other methods in single scan point recovery, where each position represents an individual amplitude/phase reconstruction from a corresponding diffraction pattern.
+This improvement is particularly evident in the vertical edge on the left side of Fig.
+3(a).
 The reconstructed images also showed contrast levels closer to the ground truth, with more natural transitions between dark and bright areas.
-By contrast, the CNN-based models tended to smooth these sharp features and introducing artifacts and distortions.
+By contrast, the CNN-based models tended to smooth these sharp features and introducing artifacts and distortions accuracy.
+Figure 4: Frequency analysis comparison of full-scene ptychographic retrievals on simulated data.
+(a) 1D diagonal cross-sections of average 2D PSD from stitched simulations.
+Red curve: ground truth; Blue curve: models.
+The red box represents obvious abnormal model retrieval.
+(b) Quantitative breakdown of energy distribution across low, mid, and high frequency bands.
+Frequency ranges are defined based on the radial distance from the PSD center, with boundaries at 1/3 and 2/3 of the maximum frequency.
 IV-B2 Full-stitched Field Retrieval Analysis
-Table I presents full-field stitching results under ideal (noise-free) and realistic (noisy) conditions, demonstrating Ptychoformer’s superior performance across all metrics.
-In ideal conditions, Ptychoformer achieved 16.0% lower MSEamp and 6.7% higher PSNRamp compared to PtychoNN, with similar improvements under noisy conditions.
+Table I presents full-field stitching results under ideal (noise-free) and realistic (noisy) conditions, demonstrating PPN’s superior performance across all metrics.
+In ideal conditions, PPN achieved 16.0% lower MSEamp and 6.7% higher PSNRamp compared to PtychoNN, with similar improvements under noisy conditions.
 Fig.
-4(c) visually confirms Ptychoformer’s superior global consistency and minimal boundary artifacts, particularly in reconstructing vertical column ends and background textures.
+3(c) visually confirms PPN’s superior global consistency and minimal boundary artifacts, particularly in reconstructing vertical column ends and background textures.
 The dual-branch structure enables our model to effectively capture both low-frequency and mid-to-high frequency information of diffraction patterns.
 This capability stems from the model’s ability to capture the intrinsic geometric structure of the data, analogous to continuous mapping on high-dimensional manifolds.
-Repeated measures ANOVA and pairwise $t$-tests (Bonferroni correction, $\alpha=0.0083$) showed Ptychoformer significantly outperformed the other methods in all metrics in both conditions ($p<0.0083$), especially in MSEamp and SSIMamp ($p<0.0001$).
+Repeated measures ANOVA and pairwise $t$-tests (Bonferroni correction, $\alpha=0.0083$) showed PPN significantly outperformed the other methods in all metrics in both conditions ($p<0.0083$), especially in MSEamp and SSIMamp ($p<0.0001$).
 The three CNN-based methods showed no significant differences, performing similarly across most metrics.
+Fig.
+3 focuses on intra-family comparisons among deep learning architectures to isolate the impact of structural variations under consistent training supervision.
+Comparative evaluations with conventional iterative algorithms (e.g., ePIE) are presented separately in Section IV-E3, where differences in computational efficiency and resilience to varying overlap ratios are quantitatively assessed and visually demonstrated.
+IV-B3 Spatial Frequency Fidelity Analysis
+We analyze frequency-domain fidelity using 1D diagonal cross-sections of the averaged power spectral density (PSD), defining three bands by normalized radial distance: low (0-1/3 Nyquist), mid (1/3-2/3), and high (2/3-1) frequencies.
+[…]
 IV-C Validation on Experimental Synchrotron Data
 Figure 5: Performance comparison on real experimental samples.
-(a,b) Visual comparison of retrieved amplitude and phase.
-For a specific hook-shaped detail, only our model effectively restores it, significantly outperforming CNN-based methods in fine structure retrieval.
+(a,b) Visual comparison of retrieved amplitude and phase (scale bar: 200nm).
+For a specific hook-shaped detail in ROI (region of interest), only our model effectively restores it, significantly outperforming CNN-based methods in fine structure retrieval.
 (c-e) Quantitative comparison of PSNR, SSIM, and MSE.
+Figure 6: Comparison of generalization capabilities across different methods using the Fudan University logo (simulated data) as a test sample with distribution significantly different from the training set.
+(a) Phase retrieval results with intensity values normalized to $\pi$.
+(b) Detailed ROI comparison.
 Fig.
-5 demonstrates Ptychoformer’s superior performance on real experimental data, consistent with our findings from simulated datasets.
-Across all metrics (PSNR, SSIM, and MSE) for both amplitude and phase retrieval, Ptychoformer consistently outperforms other models.
+5 demonstrates PPN’s superior performance on real experimental data, consistent with our findings from simulated datasets.
+Across all metrics (PSNR, SSIM, and MSE) for both amplitude and phase retrieval, PPN consistently outperforms other models.
 These improvements are visible in fine structural details (Fig.
-5a,b), indicating Ptychoformer’s enhanced resolution and fidelity in practical imaging applications.
+5 (a,b)).
+The model’s ability to preserve edge sharpness and resolve intricate features in experimental data validates its potential for practical applications where accurate reconstruction of complex nanostructures is essential for scientific interpretation.
 IV-D Generalization Capabilities Comparison
-We evaluated Ptychoformer’s generalization by trained on simulated tungsten patterns with straight-line features and evaluated on complex, curved patterns without fine-tuning, assessing the model’s ability to extrapolate learned features to significantly different sample geometries.
+We evaluated PPN’s generalization by training on simulated 1951 USAF pattern (the same with Section IV-B) with straight-line features and evaluating on complex, curved patterns (Fig.
+6), without fine-tuning, assessing the model’s ability to extrapolate learned features to significantly different sample geometries.
+This evaluation connects directly to practical applications where pre-trained models must process new samples with unknown structures—a scenario where acquiring specific training data is often impractical.
+In these contexts, generalization capability determines the practical utility of deep learning for ptychographic imaging.
 Fig.
-6 shows Ptychoformer’s superior retrieval of curved patterns.
-Quantitative results in Table II support these visual observations.
-For amplitude retrieval, Ptychoformer achieves a PSNRamp of 8.92 dB and SSIMamp of 59.06%, outperforming other models by 0.83-0.95 dB and 8.51-11.28%, respectively.
-For phase retrieval, Ptychoformer attains a PSNRphase of 8.68 dB and SSIMphase of 49.44%, surpassing other models by 0.91-1.04 dB and 32.16-35.71%, respectively.
-These results demonstrate Ptychoformer’s ability to capture fundamental relationships between diffraction patterns and sample structures, aligning with the geometry-agnostic nature of ptychographic phase retrieval.
-Figure 6: Comparison of generalization capabilities across different models: Direct test on a new unseen sample with significantly different distribution from the training set.
-(a) Amplitude and (b) phase retrieval of a single diffraction pattern.
-(c) Amplitude and (d) phase retrieval of a full scene (Fudan University logo).
-Models compared: Ground truth, PtychoNN, PtyNet, DPI, and Ptychoformer.
-TABLE II: Performance Comparison for Out-of-Distribution Generalization.
-Model Amplitude Phase
-PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
-PtychoNN 8.03 50.55 7.77 37.34
-PtyNet 7.97 47.78 7.64 36.43
-DPI 8.09 50.14 7.65 37.41
-Ptychoformer 8.92 59.06 8.68 49.44
+6 compares phase retrieval results across different methods.
+While deep learning models cannot match the reconstruction quality of the ePIE algorithm, PPN outperforms all other DL-based methods.
+The red boxes highlight high-frequency details where PPN preserves spatial information with minimal artifacts compared to other end-to-end methods.
+Results show that even with different feature distributions between training and testing data, PPN achieves superior reconstruction quality and better preservation of high-frequency spatial details.
 IV-E Performance Comparison Under Limited Data
@@
 This is particularly problematic for radiation-sensitive materials [31] and in situ dynamic studies [32].
-Our study focuses on this aspect, evaluating model performance under two kinds of limited data conditions.
-IV-E1 Impact of Training Set Size
-Figure 7: Training and testing set divisions of different sizes. Black boxes represent training set sizes, while white boxes represent testing sets, shown as squares.
+Here, the overlap ratio is defined as the physical overlap between adjacent scanning positions in real space.
+Figure 7: Model performance evaluation on simulated data.
+(a) Reconstruction quality versus training data size at overlap ratio 75%.
+(b) Performance stability under different test set overlap ratios with fixed training set size (100 samples).
+IV-E1 Comparison with Different Training Data Size
+We evaluated the effect of training data size on model performance by training models with 10 different random initializations while keeping all hyperparameters consistent and testing on the same test set as in Section IV-B.
+As shown in Fig.
+7(a), when the training samples decrease from 2928 to 50, PPN demonstrates superior data efficiency: the PSNR for amplitude reconstruction decreases by only 35%, and phase reconstruction by 15%.
+In contrast, PtychoNN shows the most significant performance degradation, with PSNR dropping by approximately 70% for both amplitude and phase reconstruction.
+This resilience, analyzable through compressed sensing theory.
+[33], suggests PPN learns an optimal sparse prior.
+IV-E2 Comparison Across Overlap Ratios with a Fixed Training Size
+We investigated the impact of test set overlap ratios (0%, 25%, 50%, and 75%) while maintaining a fixed training set size of 100 samples, consistent with the models used in Section IV-E1.
+While the test area maintains consistent spatial coverage (matching Fig.
+3 (c)’s field of view), the number of test samples decreases with lower overlap ratios due to reduced spatial sampling density.
+As illustrated in Fig.
+7(b), all models exhibit remarkable metric stability across different overlap ratios.
+Taking PPN as an example, its amplitude reconstruction PSNR remains stable between 9.5–9.6 ($\Delta=0.1$) and SSIM between 0.39–0.40 ($\Delta=0.01$), with one-way ANOVA ($p>0.05$) confirming no significant differences between overlap ratio groups.
+This overlap-invariant behavior stems from our training approach, which uses simulated real-space amplitude and phase as ground truth—data that typically requires exceptionally high overlap ratios to obtain in conventional experimental workflows.
+This characteristic enables an optimal workflow: using high-overlap data for model training while deploying at low overlap during testing.
+These findings naturally lead to the question of how PPN compares to traditional iterative methods across different overlap ratios, which we explore in the following section.
+Figure 8: Comparison between ietrative method and PPN across different overlap ratios.
+(a) SSIM and reconstruction time comparison.
+(b) and (c) Reconstructed images using iterative method and our proposed PPN, respectively, at different overlap ratios (30%, 60%, 90%).
+Figure 9: Ablation study performed on simulated data showing amplitude and phase reconstructions with 2D images and 1D cross-sectional profiles , comparing ground truth (blue solid lines) with AI inference results (red dashed lines).
+IV-E3 Comparison with Iterative Methods Across Overlap Ratios
+We compared PPN with the iterative method (where we use the widely-used ePIE as our baseline) across various overlap ratios to assess both reconstruction quality and computational efficiency.
 Fig.
-8 (a-c) illustrates Ptychoformer’s performance under varying training set sizes.
-As the training size decreases from 2928 to 50, Ptychoformer consistently outperforms CNN-based models, especially at smaller training set sizes.
-Comparing with DPI, the best-performing CNN method, we showcase results at representative training sizes of 1000 and 50.
-Even with 1000 samples, DPI exhibits limitations with complex structures like the top-right crosshatch pattern, producing block-shaped high-frequency artifacts.
-Ptychoformer, in contrast, maintains superior resolution.
-At the minimal training size of 50 samples, DPI’s retrieval quality significantly deteriorates, while Ptychoformer still distinguishes basic structural features like crosshatches and L-shaped patterns.
-This robustness is attributed to Ptychoformer’s PCA mechanism and dual-branch architecture, which align with diffraction physics and capture multi-scale features effectively, demonstrating superior ability to learn from limited data.
-IV-E2 Performance Across Overlap Ratios
-Fig.
-8 (d) demonstrates Ptychoformer’s effectiveness at varying overlap ratios.
-At 0% overlap, Ptychoformer achieves higher PSNRamp, surpassing CNNs.
-From 75% to 0% overlap, Ptychoformer’s PSNRamp drops by a smaller amount compared to CNNs.
-This resilience, analyzable through compressed sensing theory.
-[32], suggests Ptychoformer learns an optimal sparse prior.
-Lower overlap requirements could enable faster experiments with reduced radiation damage.
-Ptychoformer’s efficiency demonstrates the synergy between deep learning and domain-specific knowledge in ptychography.
-Figure 8: Performance comparison across training set sizes and overlap ratios.
-(a,b) Retrievals with 1000 and 50 training samples.
-(c,d) PSNR and SSIM metrics for varying training sizes and overlap ratios.
-Higher values for both PSNR and SSIM indicate better performance.
-TABLE III: Ablation Study Results
+8 presents this comparison, with PPN trained using the same dataset described in Section IV-B, where sample features remain within the learned distribution.
+As shown in Fig.
+8(a), PPN maintains SSIM values between 0.86–0.92 across all overlap ratios, indicating consistent reconstruction quality.
+In contrast, iterative method exhibits strong overlap dependence, with SSIM dropping from 0.95 at 90% overlap to only 0.12 at 30% overlap.
+At 30% overlap, PPN completes reconstructions in approximately 0.15 seconds versus iterative method’s 125 seconds, representing an 852$\times$ speed improvement, and at 60% overlap, a 1767× speed improvement.
+The reconstruction time was measured by averaging over ten consecutive tests, with error bars representing the variance.
+Reducing overlap from 90% to 30% increases data acquisition efficiency by approximately 49 times for a fixed area, as scan point density is inversely proportional to the square of the step size.
+Theoretically, this combined effect could yield up to 41,748$\times$ (49 $\times$ 852) overall efficiency improvement.
+This enables significantly faster experiments with substantially reduced radiation exposure while still producing high-quality reconstructions.
+This separation of training and testing strategies, along with PPN’s efficiency, demonstrates the synergy between deep learning and domain-specific knowledge in ptychography, providing a viable technical pathway for achieving high-throughput imaging.
+IV-F Ablation Study
+TABLE II: Ablation Study Results on Simulated Data
 Model Variant Amplitude Phase
@@
 Ours w/o D 11.02 60.50 17.32 54.50
-Ours w/o OC 14.21 86.10 18.39 69.40
 Ours 14.42 87.00 19.41 70.60
-IV-F Ablation Study
-To justify each component of the proposed Ptychoformer, we conducted comprehensive ablation studies.
-We compared several variants of our model: 1) Ours without NonLocal Coherence Branch (Ours w/o NLB), using only the Local Dependencies Branch; 2) Ours with Multi-Head Self-Attention (MHSA) in NLB (Ours w/ MinN), replacing PCA with standard MHSA in the NonLocal Branch; 3) Ours without Decoder (Ours w/o D), using fully connected layers instead of CNN for upsampling in the decoder; 4) Ours without Output Constraints (Ours w/o OC), removing the amplitude and phase constraints in the output layer; and 5) Ours (Full Model), the complete proposed Ptychoformer model.
-Table III presents the quantitative results of these variants, with the best performance in bold.
-The NonLocal Coherence Branch, reflecting diffraction patterns’ global nature, showed a modest impact, primarily affecting structural integrity.
-PCA demonstrated clear advantages over standard MHSA, with its replacement by MHSA causing a 4.16% drop in PSNRamp and 3.56% in SSIMamp.
-The CNN decoder proved crucial, its absence leading to a substantial 23.58% decrease in PSNRamp and 30.46% in SSIMamp.
-This significant impact is attributed not only to its role in maintaining spatial coherence during upsampling but also to its inductive bias, which enables effective learning from limited data.
-Output constraints, while showing a modest quantitative impact, play a vital role in enforcing physical consistency.
+To justify each component of the proposed PPN, we conducted comprehensive ablation studies.
+As shown in Fig.
+9, we compared both amplitude and phase reconstructions, including 2D reconstructed images and 1D cross-sectional profiles.
+We examined several model variants: 1) Ours without NonLocal Coherence Branch (Ours w/o NLB), which means using only the Local Dependencies Branch and removing the NonLocal Coherence Branch that contains proposed PoCA; 2) Ours with Multi-Head Self-Attention (MHSA) in NLB (Ours w/ MinN), replacing PoCA with standard MHSA in the NonLocal Branch; 3) Ours without Decoder (Ours w/o D), using fully connected layers instead of CNN for upsampling in the decoder; and 4) Ours (Full Model), the full proposed model.
+The ablation experiments were conducted using the same training set, test set, and parameter settings as in Section IV-B.
+The quantitative results are presented in Table II.
+The removal of the NonLocal Coherence Branch leads to decreased structural coherence in phase reconstruction, which is clearly observable in the 1D profiles.
+Replacing PoCA with standard MHSA results in a 4.16% drop in PSNRamp and 3.56% in SSIMamp, with poorer edge transitions visible in the profile plots.
+The absence of the CNN decoder causes the most significant performance degradation, with PSNRamp and SSIMamp decreasing by 23.58% and 30.46%, respectively.
+The reconstructed images without the CNN decoder appear blurry, consistent with the experimental results in [17].
+The 1D profiles reveal that without the CNN decoder, the model struggles to maintain accurate amplitude levels and phase transitions.
+The full model achieves the best performance across all metrics, particularly in regions with dramatic phase changes in the 1D cross-sectional profiles.
 These results validate the effectiveness of each proposed component in our architecture.
+IV-G Model Variants
+To validate our model’s effectiveness, we benchmarked against two widely-adopted hybrid CNN-ViT architectures from computer vision: TransUNet [34] and SegFormer [35].
+These baselines remain relevant in recent research across multiple domains [36, 37, 38, 39], demonstrating their continued utility for important results in various fields.
+The first variant adopts a TransUNet-inspired architecture [34], employing a hybrid CNN-Transformer encoder where CNNs extract initial features before Transformer processing, with a decoder utilizing a cascading structure and skip connections.
+The second implements a SegFormer-based approach [35], featuring a hierarchical structure with progressive resolution reduction for multi-scale feature extraction and replacing the first linear layer in the feed-forward network with a $3\times 3$ convolution, similar to the approach used in PtychoFormer[26].
+PtychoFormer uses an encoder based on SegFormer and a decoder adapted specifically for ptychography tasks.
+Experiments conducted under noise-free conditions (see Section IV-B2) demonstrate that PPN outperforms all baseline models (Table III).
+The PtychoFormer results are comparable to our SegFormer-inspired model due to their similar structures.
+These variants’ limitations stem from designs optimized for real-space image processing: The TransUNet variant, while benefiting from the combination of CNN and Transformer, still relies on local feature extraction in its initial stages.
+This approach is suboptimal for capturing the global coherence information present in diffraction patterns.
+Similarly, the SegFormer variant’s hierarchical structure, while effective for multi-scale feature extraction in natural images, may lose critical high-frequency information in the context of diffraction patterns due to its progressive downsampling.
+TABLE III: Performance comparison of Model Variants.
+Method Amplitude Phase
+PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
+TransUNet[34] 13.57 81.20 18.77 67.80
+SegFormer[35] 13.76 83.30 19.12 68.10
+PtychoFormer[26] 13.77 83.54 19.17 68.18
+PPN (Ours) 14.42 87.00 19.41 70.60
 V Discussion
-V-A Frequency Analysis
-Further analysis of the noise-free full-stitched field using 1D diagonal cross-sections of the average 2D power spectral density (PSD) and frequency distribution statistics (Fig.
-9).
-In Fig.
-9(a), Ptychoformer’s PSD curve closely follows the ground truth, while the DPI model exhibits significant mid and high-frequency fluctuations, indicating potential artifacts.
-Quantitatively, Ptychoformer achieves a frequency distribution (99.75%, 0.22%, 0.03% for low, mid, and high frequencies, respectively), closely approximating the ground truth (99.58%, 0.38%, 0.04%).
-This performance exceeds that of CNN-based models, which show mid and high frequency degradation (e.g., PtychoNN: 99.93%, 0.06%, 0.01%).
-These results highlight Ptychoformer’s efficacy in extracting mid-to-high frequency information and retrieving corresponding detailed features, leading to superior artifact suppression.
-Figure 9: Frequency analysis comparison of full-scene ptychographic retrievals.
-(a) 1D diagonal cross-sections of average 2D PSD from stitched simulations.
-Red curve: ground truth; Blue curve: models.
-The red box represents obvious abnormal model retrieval.
-(b) Quantitative breakdown of energy distribution across low, mid, and high frequency bands.
-Frequency ranges are defined based on the radial distance from the PSD center, with boundaries at 1/3 and 2/3 of the maximum frequency.
-Figure 10: Performance analysis of weighted loss function $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ in ptychographic retrieval.
+Figure 10: Comparison of different models in terms of their complexity (parameters and FLOPS) (left) and training efficiency (convergence epochs and time per epoch) (right).
+V-A Computational Complexity Analysis
+Our analysis compares PPN with CNN-based methods (PtychoNN, PtyNet, DPI) and the ViT-CNN hybrid baseline (PtychoFormer[26]) through computational metrics in Fig.
+10.
+While maintaining hierarchical CNN decoders common to all compared methods, our encoder design achieves substantial efficiency gains.
+Among CNN approaches, PPN demonstrates intermediate complexity (6.12 MB parameters) between PtyNet (1.24 MB) and DPI (57.14 MB), yet converges faster (30 epochs) than both PtychoNN (45 epochs) and DPI (35 epochs).
+The 0.382 GFLOPs cost represents a 63.7% reduction from DPI’s 1.049 GFLOPs, confirming efficient feature learning despite transformer integration.
+Compared to PtychoFormer’s ViT-CNN architecture, our model reduces parameters by 91.1% (6.12 MB vs.
+68.96 MB) and FLOPs by 65.4% (0.382 G vs.
+1.105 G) while matching convergence speed (30 epochs).
+This efficiency comes from three optimizations: 1) fixed feature dimensions preventing channel inflation, 2) parallel dual-branch processing instead of hierarchical refinement, and 3) physics-inspired attention limiting learnable parameters.
+These changes eliminate redundant operations while preserving physical priors.
+Figure 11: Performance analysis of weighted loss function $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ on simulated data.
 PSNR, SSIM are plotted against $\alpha$ for amplitude and phase retrievals.
 V-B Loss Function Analysis
-To optimize retrieval quality in ptychographic phase retrieval, we evaluated various loss functions (as shown in Table IV.
-Standard metrics like MSE and Mean Absolute Error (MAE) showed similar performance, while the Huber loss [33], defined as: $L_{\text{Huber}}(x)=\begin{cases}\frac{1}{2}x^{2}&\text{for }|x|\leq\delta\\ \delta(|x|-\frac{1}{2}\delta)&\text{otherwise}\end{cases}$ offered robustness to outliers.
-The Negative Pearson Correlation Coefficient (NPCC) [34] loss, despite its success in single-frame phase retrieval [12], showed poor convergence stability in our experiments.
-This is likely due to its sensitivity to local statistics in small ptychographic patches, contrasting with its effectiveness on larger, more statistically stable single-frame images.
-We introduced a novel combined loss function inspired by the work on image restoration with neural networks [35]: $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ where $L_{\text{NSSIM}}=-\text{SSIM}(I_{q},|\mathcal{F}{P(r)\psi(r)}|^{2})$, where $\psi(r)$ defined as $\psi(r)=P(r-r_{j})O(r)$.
-This approach balances global consistency (MSE) with local structural preservation ( Negative Structural Similarity Index Measure (NSSIM) [36]).
-Experiments revealed optimal performance at $\alpha=0.9$, with consistent improvement as $\alpha$ increased from 0.2 to 0.9, as shown in Fig.
-10.
-This robustness to $\alpha$ values reduces the need for precise tuning in practical applications.
-The combined loss function excels in addressing affine ambiguity, as $\nabla L_{\text{MSE}}$ is sensitive to global affine transformations $T$ ($\frac{\partial L_{\text{MSE}}}{\partial T}\neq 0$), while $\nabla L_{\text{NSSIM}}$ maintains local structure invariance ($\frac{\partial L_{\text{NSSIM}}}{\partial(\text{local structure})}\approx 0$), forming a more robust optimization target.
 TABLE IV: Comparison with Different Loss Functions
@@
 Note:The weighted loss combines MSE and NSSIM losses, where $\alpha$ represents the weight of MSE (0.9) and 1-$\alpha$ (0.1) is the weight of NSSIM.
-V-C Model Variants
-To validate the effectiveness of our physics-informed design, we compared Ptychoformer with two state-of-the-art hybrid CNN-ViT models, adapted for ptychography task.
-The first is a TransUNet-inspired model [37], which uses a hybrid CNN-Transformer encoder where CNN extracts features and the Transformer processes them further.
-Its decoder employs a cascading structure with skip connections.
-The second is a SegFormer-based model [38], which features a hierarchical structure with progressive resolution reduction for multi-scale features.
-This model replaces the first linear layer in the feed-forward network with a $3\times 3$ convolution.
-We conducted the same experiment as in Section IV-B2 under noise-free conditions.
-Table V shows that Ptychoformer outperforms both models.
-TransUNet and SegFormer’s limitations stem from designs optimized for real-space image processing.
-The TransUNet variant, while benefiting from the combination of CNN and Transformer, still relies on local feature extraction in its initial stages.
-This approach is suboptimal for capturing the global coherence information present in diffraction patterns.
-Similarly, the SegFormer variant’s hierarchical structure, while effective for multi-scale feature extraction in natural images, may lose critical high-frequency information in the context of diffraction patterns due to its progressive downsampling.
-TABLE V: Performance comparison of Model Variants.
-Method Amplitude Phase
-PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
-TransUNet[37] 13.57 81.20 18.77 67.80
-SegFormer[38] 13.76 83.30 19.12 68.10
-Ptychoformer 14.42 87.00 19.41 70.60
+To optimize retrieval quality in ptychographic phase retrieval, we evaluated various loss functions (as shown in Table IV.
+Standard metrics like MSE and Mean Absolute Error (MAE) showed similar performance, while the Huber loss [40], defined as: $L_{\text{Huber}}(x)=\begin{cases}\frac{1}{2}x^{2}&\text{for }|x|\leq\delta\\ \delta(|x|-\frac{1}{2}\delta)&\text{otherwise}\end{cases}$ offered robustness to outliers.
+The Negative Pearson Correlation Coefficient (NPCC) [41] loss, despite its success in single-frame phase retrieval [12], showed poor convergence stability in our experiments.
+This is likely due to its sensitivity to local statistics in small ptychographic patches, contrasting with its effectiveness on larger, more statistically stable single-frame images.
+We introduced a novel combined loss function inspired by the work on image restoration with neural networks [42]: $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ where $L_{\text{NSSIM}}=-\text{SSIM}(I_{q},|\mathcal{F}{P(r)\psi(r)}|^{2})$, where $\psi(r)$ defined as $\psi(r)=P(r-r_{j})O(r)$.
+This approach balances global consistency (MSE) with local structural preservation ( Negative Structural Similarity Index Measure (NSSIM) [43]).
+Experiments revealed optimal performance at $\alpha=0.9$, with consistent improvement as $\alpha$ increased from 0.2 to 0.9, as shown in Fig.
+11.
+This robustness to $\alpha$ values reduces the need for precise tuning in practical applications.
+The combined loss function excels in addressing affine ambiguity, as $\nabla L_{\text{MSE}}$ is sensitive to global affine transformations $T$ ($\frac{\partial L_{\text{MSE}}}{\partial T}\neq 0$), while $\nabla L_{\text{NSSIM}}$ maintains local structure invariance ($\frac{\partial L_{\text{NSSIM}}}{\partial(\text{local structure})}\approx 0$), forming a more robust optimization target.
 VI Conclusion
-In this study, we introduced Ptychoformer, a deep learning framework designed for ptychographic retrieval from frequency-domain data.
-By incorporating physical insights into the model architecture, particularly through the PCA mechanism and dual-branch structure, Ptychoformer demonstrated superior performance over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency.
-The principles underlying Ptychoformer show promise for extension to other frequency-domain inverse imaging problems, potentially advancing high-resolution retrieval techniques across various scientific disciplines.
+We propose PPN, a physics-inspired deep learning framework that significantly improves ptychographic reconstruction through two core innovations: (1) an architecture combining local feature extraction with global pattern coherence modeling, and (2) a novel attention mechanism tailored for diffraction physics.
+PPN demonstrated superior performance over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency.
+Current limitations include performance validation primarily on simulated and small-scale experimental data – future work will focus on large-scale real-world deployments across diverse imaging conditions.
+The framework’s physics-inspired design principles show promising extensibility to other frequency-domain inverse problems like cryo-EM and astronomical imaging.
 Acknowledgments